### v3

# Very Basic Set theory

Very Basic Set theory
Doeko H. Homan
(Date: August 18, 2024)
George Boolos describes the naive conception of set: ”Any predicate has an extension” (Boolos [1971] pp.215-231). Then Russell’s paradox is used to prove the inconsistency of naive set theory. Russell’s paradox reads:
”No set can contain all and only those sets which do not contain themselves”.
We call a set which contains itself an ‘individual’.
In this article we show Russell’s paradox is logically equivalent to:
For every set $s$, an individual is contained in $s$ or a set that is not an individual
is not contained in $s$.
Therefore Very basic set theory is a set theory with individuals.
1. What are individuals?
In (Kanamori [2003] pp.285-288) Akihiro Kanamori describes that with the ‘empty
extension’ and the ‘axiom of restriction’ there is no need for individuals (also
known as atoms or urelemente). Individuals ”are objects distinct from the
null set yet having no members and capable of belonging to sets”.
In (Boolos [1971] pp.216-217) ”individuals (= non-sets)”.
In section 4.6.5 (Mendelson [2015]) Elliott Mendelson explains:
”The words ‘individual’ and ‘atom’ are sometimes used as synonyms for
‘urelement’. Thus, urelements have no members”. And in a footnote (p.304):
”Zermelo’s 1908 axiomatization permitted urelements”.
In (Zermelo [1908] pp.262-265) Ernst Zermelo describes in ”10. Theorem.”:
”Für jedes Element $x$ von $M$ ist es definit, ob $x$ $\varepsilon$ $x$ oder nicht; diese Möglichkeit
$x$ $\varepsilon$ $x$ ist an und für sich durch unsere Axiome nicht ausgeschlossen.”
Thus ‘$x$ $\varepsilon$ $x$ or not’ is definite; the possibility ‘$x$ $\varepsilon$ $x$’ is not excluded by our axioms.
Therefore Zermelo’s 1908 axiomatization permitted sets which contain themselves.
Sets are denoted by $s$, $t$, $u$, …. Symbol $\in$ is the ‘membership relation’ symbol. Then $s\in t$ means $t$ contains $s$ or $s$ is a member of $t$. The negation of $s\in t$ is denoted by $s\notin t$. The membership relation is definite, that is for every $s$ and for every $t$ ($s\in t$ or $s\notin t$) but never ($s\in t$ and $s\notin t$). If $s\in s$ then $s$ is an individual.
We prove Russell’s paradox.
Assume exists $s$ such that for every $u$ ($u\in s$ if and only if $u\notin u$).
The formula applies to every $u$ thus we can choose $s$ therefore
exists $s$ such that ($s\in s$ if and only if $s\notin s$).
The question ‘is $s$ contained in $s$’ or ‘is $s$ not contained in $s$’ is definite but there is
no answer: If $s\in s$ then $s\notin s$ and vice versa. Thus $s$ is nonexisting. Therefore:
”No set can contain all and only those sets which do not contain themselves”. $\Box$
The membership relation is definite thus for every $s$ ($s\in s$ or $s\notin s$).
If $s\in s$ then exists $u$ such that ($u\in s$ and $u\in u$),
if $s\notin s$ then exists $u$ such that ($u\notin s$ and $u\notin u$) therefore:
For every $s$ exists $u$ such that (($u\in s$ and $u\in u$) or ($u\notin s$ and $u\notin u$)).
That is logically equivalent to Russell’s paradox:
It is not the case that exists $s$ such that for every $u$ ($u\in s$ if and only if $u\notin u$).
Thus for every $s$, an individual is contained in $s$ or a set that is not an individual
is not contained in $s$. $\Box$
Zermelo’s 1908 axiomatization permitted individuals. And the prove of Russell’s paradox includes individuals and other sets. But a set which contains itself, an individual, is problematic. (Boolos [1971] p.217, pp.219-220):
”perhaps the mind ought to boggle at the idea of something’s containing itself.”
”It is important to realize how odd the idea of something’s containing itself is.”
”Here are some things. Now we bind them up into a whole. Now we have a set.”
However, ‘binding up some things’ or ‘containing something’ is not the definition of a set. (Boolos [1971] p.215):
”Cantor also defined a set as a ‘many, which can be thought of as one, i.e. a
totality of definite elements that can be combined into a whole by a law.’ ”
Then ”which can be thought of as one,” a set is a ‘mental picture’ in your mind. Therefore we prefer to write ‘constituent of a set’ rather than ‘contained in a set’.
Set theory is a language that follows the rules of logic to communicate the mental picture of a set with the physical world. Logic treats every possibility. Individuals are possible. Then, hopefully, someone reading or listening that language forms a correct mental picture in her/his mind. That is called ‘experience’.
What is an individual? ‘Set $s$ is an individual’ is what is called in (Wittgenstein [1999]) ”an atomic fact”. In the Introduction (p.12) Bertrand Russell explains:
”If an atomic fact is analyzed as fully as possible (theoretical, not practical
possibility is meant) the constituents finally reached may be called simples
or objects.”
”(…) the naming of simples is shown to be what is logically first in logic.”
Therefore in Very basic set theory the naming of individuals is logically first.
2. What is a set?
At an early age you develop the mental picture of ‘what is a set’. You belong to a family or a clan, to the inhabitants of a village or a particular region. Then experience shows there are things that constitute a set. The members of a set are the constituents of that set. There is a membership relation between sets.
Very basic set theory is based on
the world of individuals,
the membership relation is definite,
the basic properties of a set are the constituents of that set and the sets that set is a member of,
the only constituent of an individual is that individual itself.
Experience shows a set ‘is equal to’ or ‘is not equal to’ another set. Sets with different constituents are different sets, and different sets have different constituents.
3. Very basic set theory
A ‘formula’ is a description of properties of sets. Thus $s\in t$ is a formula. Formulas are denoted by $\Phi$, $\Psi$, $\ldots$ and are built up from formulas by means of negation $\neg\Phi$ (‘not $\Phi$’) and conjunction $\Phi\wedge\Psi$ (‘$\Phi$ and $\Psi$’). Negation has higher precedence than conjunction. The parenthesis (((( and )))) preclude ambiguity and improve legibility. Then $s\notin t$ is short for the formula $\neg(s\in t)$.
If $\Phi$ is a formula then $(\Phi)$, $\forall s(\Phi)$ and $\exists s(\Phi)$ are formulas. $\forall s(\Phi)$ means ‘for every set $s$ $\Phi$’, and $\exists s(\Phi)$ means ‘exists set $s$ such that $\Phi$’. Then e.g. $\neg\exists s\forall u(\Phi)$ is short for the formula $\neg(\exists s(\forall u(\Phi)))$.
If there is $s$ in $\Phi$ not guided in $\Phi$ by $\forall s$ or by $\exists s$ then ‘$s$ is free in $\Phi$’. If $s$ is free in $\Phi$ and $\neg\exists s(\Phi\wedge\neg\Phi)$ then ‘$\Phi$ is definite’.
We define the logical connectives $\vee$, $\rightarrow$ and $\leftrightarrow$. The precedence of each of it is lower than conjunction.
$\Phi\vee\Psi$ (‘$\Phi$ or $\Psi$ or both’) is short for the formula ¬($\neg(\neg\Phi\wedge\neg\Psi)$,
$\Phi\rightarrow\Psi$ (‘$\Phi$ implies $\Psi$’) is short for the formula ($\neg\Phi\vee\Psi$),
$\Phi\leftrightarrow\Psi$ (‘$\Phi$ if and only if $\Psi$’) is short for the formula ($(\Phi\rightarrow\Psi)\wedge(\Psi\rightarrow\Phi))$.
If $\Phi$ and $\Psi$ are definite then $\neg\Phi$, $\Phi\wedge\Psi$, $\Phi\vee\Psi$, $\Phi\rightarrow\Psi$ and $\Phi\leftrightarrow\Psi$ are definite.
Then Russell’s paradox is the formula $\neg\exists s\forall u(u\in s\leftrightarrow u\notin u)$, and is logically equivalent to $\forall s\exists u((u\in s\wedge u\in u)\vee(u\notin s\wedge u\notin u))$.
The naive conception of set is inconsistent: formula $\exists s\forall u(u\in s\leftrightarrow u\notin u)$ is not definite. Therefore we postulate ‘specification’ (also known as ‘separation’).
  • if $u$ is free in $\Phi$ and $\Phi$ is definite then $(u\in s\wedge\Phi)$ is definite and
    (specification) $\forall s(\exists u(u\in s\wedge\Phi)\rightarrow\exists v\forall u(u\in v\leftrightarrow(u\in s\wedge\Phi)))$.
Thus the conjunction of mutually exclusive properties does not constitute a set.
We prove: If $\exists u(u\in s\wedge u\in u)$ then exists $v$ whose members are all and only
those individuals that are constituent of $s$.
Formula $(u\in s\wedge u\in u)$ is definite and is not the conjunction of mutually exclusive properties. Apply ‘specification’ to $s$ to find $v$
$\exists v\forall u(u\in v\leftrightarrow(u\in s\wedge u\in u))$. Then an individual is a constituent of $v$.
The formula applies to every $u$ thus we can choose $v$
$\exists v(v\in v\leftrightarrow(v\in s\wedge v\in v))$ therefore
$\exists v((v\notin v\vee(v\in s\wedge v\in v))\wedge((v\notin s\vee v\notin v)\vee v\in v))$.
Then $\exists v(v\notin v\vee v\in s)$. Thus $v$ is not an individual, or $v$ is an individual and a constituent of $s$. A set of all individuals is not excluded. $\Box$
We prove ‘the set of all sets’ is nonexisting.
If $\exists u(u\in s\wedge u\notin u)$ then formula $(u\in s\wedge u\notin u)$ is definite and is not the conjuction of mutually exclusive properties.
Apply ‘specification’ to $s$ to find $v$
$\exists v\forall u(u\in v\leftrightarrow(u\in s\wedge u\notin u))$. Then a set is a member of $v$.
The formula applies to every $u$ thus we can choose $v$
$\exists v(v\in v\leftrightarrow(v\in s\wedge v\notin v))$ therefore
$\exists v((v\notin v\vee(v\in s\wedge v\notin v))\wedge((v\notin s\vee v\in v)\vee v\in v))$.
Then $\exists v(v\notin v\wedge v\notin s)$. Thus $v$ is not an individual and $v$ is not a member of $s$. Therefore ‘the set of all sets’ is nonexisting. $\Box$
4. More Axioms
Experience shows a set is equal to or a set is not equal to another set.
We define equality for sets.
Definition 1.
Set $s$ ‘is equal to’ set $t$, denoted by $s=t$ if $\forall u(u\in s\leftrightarrow u\in t)$.
Set $s$ ‘is not equal to’ set $t$, denoted by $s\neq t$ if $\neg(s=t)$. $\Box$
The axioms for equality are
  • (equality) $\forall s\forall t(s=t\rightarrow\forall u(s\in u\leftrightarrow t\in u))$,
  • (individuals) $\forall s(s\in s\rightarrow\forall u(u\neq s\rightarrow u\notin s))$.
Thus if $s=t$ then $s$ and $t$ have the same constituents and $s$ and $t$ are members of the same sets. And the only constituent of an individual is that individual itself.
Then $\forall s\forall t((s\in s\wedge t\notin t)\rightarrow(s\neq t\wedge t\notin s))$.
Any pair of sets constitute a set. We postulate ‘pairs’.
  • (pairs) $\forall s\forall t\exists v\forall u(u\in v\leftrightarrow(u=s\vee u=t))$.
Set $v$ is denoted by $\{s,t\}$ thus $\{s,t\}=\{t,s\}$. The ‘singleton’ $\{s\}$ of $s$ is $\{s,s\}$.
$s\notin s\rightarrow\{s\}\neq s$ otherwise $s\in s$. Therefore $s\notin s\rightarrow\{s\}\notin\{s\}$,
$s\in s\rightarrow\{s\}=s$, the singleton of an individual is that individual itself,
$s\neq t\rightarrow\{s,t\}\notin\{s,t\}$. Thus a pair of different sets is not an individual.
However, if $s\neq t$ and $(s\in t\wedge t\in s)$ then $s\notin s$ and $t\notin t$. Then exists $\{s,t\}$ and there is the ‘vicious membership circle’ $(s\in t\wedge t\in s)$.
The naming of individuals is logically first. Thus every set has a constituent, and vicious membership circles do not exist. Therefore we postulate ‘regularity’.
  • (regularity) $\forall s\exists v(v\in s\wedge(v\notin v\rightarrow\forall u((u\in s\wedge u\in v)\rightarrow u\in u)))$.
If $s$ is an individual or a set of individuals then $v$ is an individual. Otherwise exists $v$ and $v$ is not an individual and $s$ and $v$ only have individuals in common.
And it is not the case that exists $s$ such that
$\forall v(v\in s\wedge v\notin v\rightarrow\exists u(u\in s\wedge u\in v\wedge u\notin u))$.
Therefore any sequence of sets $s$, $t$, $u$, $v$, … such that each term is a member of the previous one $t\in s$, $u\in t$, $v\in u$ is nonexisting. Thus $(t\in s\wedge s\in t)$ is nonexisting therefore $\forall s\forall t(s\neq t\rightarrow(s\notin t\vee t\notin s))$.
The members of the members of a set constitute a set. We postulate ‘union’.
  • (union) $\forall s\exists v\forall u(u\in v\leftrightarrow\exists t(t\in s\wedge u\in t))$.
Set $v$ is denoted by $\bigcup s$. If $s$ is an individual or a set of individuals then $\bigcup s=s$. The ‘union $s\cup t$’ is defined by $s\cup t=\bigcup\{s,t\}$. Then $\bigcup\{\{s\},\{t\}\}=\{s,t\}$ and $s\cup(t\cup u)=(s\cup t)\cup u$. If $s$ and $t$ are individuals then $s\cup t=\{s,t\}$.
An ‘infinite’ sequence of different sets $s$, $\bigcup s$, $\bigcup(\bigcup s)$, … is nonexisting.
Any pair of different consecutive sets $\{w,\bigcup w\}$ implies $\exists v(v\in w\wedge v\notin v)$ otherwise $\bigcup w=w$. If $z=s\cup\bigcup s\cup\bigcup(\bigcup s)\ldots$ then
$\forall w((w\in z\wedge w\notin w)\rightarrow\exists v(v\in w\wedge v\in\bigcup z\wedge v\notin v))$.
Then $\forall v(v\in\bigcup z\rightarrow v\in z)$ thus $z$ contradicts regularity.
Therefore the sequence terminates: $\exists w(\bigcup w=w)$. Set $z$ is called ‘the transitive closure’ of $s$. Thus for every set the transitive closure exists. $\Box$
5. Conclusion
Very basic set theory is a set theory with individuals
the only constituent of an individual is that individual itself,
Russell’s paradox is logically equivalent to a formula with individuals,
the conjunction of mutually exclusive properties does not constitute a set,
the naming of individuals is logically first,
for every set the transitive closure exists.
Compared with ‘mainstream set theory’ a big difference is ‘what is logically first’. The logic of set theory allows for individuals, and Russell’s paradox is logically equivalent to a formula with individuals. Therefore we choose for ‘the naming of individuals is logically first’, not for the postulate of an ‘empty set’.
There are more differences. The naive conception of set is inconsistent thus axiom ‘specification’ is modified. The axiom ‘powerset’ is as usual but Cantor’s Theorem does not hold.
  • (powerset) $\forall s\exists v\forall u(u\in v\leftrightarrow\forall t(t\in u\rightarrow t\in s))$.
Set $v$ is denoted by ${\mathcal{P}}s$. Then ${\mathcal{P}}\{s\}=\{\{s\}\}$ thus the ‘number of members’ of $\{s\}$ is equal to the ‘number of members’ of ${\mathcal{P}}\{s\}$.
Cantor’s Theorem (Zermelo [1908] p.276) reads: For every $s$,
the ‘number of members’ of $s$ is less than the ‘number of members’ of ${\mathcal{P}}s$.
Thus Cantor’s Theorem does not hold.
The definition of equality and axiom ‘pairs’ are only meaningful in combination with sets that are logically first. Thus with axiom ‘regularity’. And ‘regularity’ is used with ‘union’ to prove that for every set the transitive closure exists.
In the last section we define ‘number sequences’ satisfying the Peano Axioms (PA) for natural numbers. The first axiom of PA reads: There is a ‘first natural number’. Any pair of individuals is a first number of a natural number sequence. Therefore ‘the set of all natural numbers’ is nonexisting. However, a set of all natural numbers with a specified first number is not excluded.
If in the world only one pair of individuals exists then indeed a set of all natural numbers with that pair of individuals as first number can postulated to exist. That is the way of ‘mainstream set theory’: There is no need for individuals, and the ‘empty set’ is postulated as ‘first natural number’.
6. Natural numbers
Definition 2.
Set $s$ is a ‘subset’ of $t$, denoted by $s\subseteq t$, if $\forall u(u\in s\rightarrow u\in t)$.
$s$ is ‘transitive’ if $\forall u(u\in s\rightarrow u\subseteq s)$, or equivalently $\bigcup s\subseteq s$. $\Box$
Every individual and every set of individuals is transitive. If $s$ is transitive then $\bigcup s$ and $s\cup\{s\}$ are transitive.
Theorem 1.
If $s$ is transitive and $\exists u(u\in s\wedge u\notin u)$ then a set of individuals is a
member of $s$: $\exists v(v\in s\wedge v\notin v\wedge\forall u(u\in v\rightarrow u\in u))$.
Proof. Apply regularity to $s$ to find $v\in s\wedge v\notin v$ such that
$\forall u((u\in s\wedge u\in v)\rightarrow u\in u)$. Set $s$ is transitive and $v\in s$ thus $\forall u(u\in v\rightarrow u\in s)$.
Therefore $\exists v(v\in s\wedge v\notin v\wedge\forall u(u\in v\rightarrow u\in u))$. $\Box$
If $u$ and $v$ are the only individuals belonging to transitive set $s$ then $\{u,v\}$ is the only possible set of individuals belonging to $s$. Therefore $\{u,v\}\in s\cup\{s\}$.
If $s$ and $t$ are transitive sets with transitive members and the only individuals belonging to $s$ or to $t$ are $u$ and $v$, thus $\forall w((w\in s\cup t\wedge w\in w)\rightarrow w\in\{u,v\})$, then the ‘law of trichotomy’ for $s$ and $t$ is easily proved and reads
$(s\notin s\wedge t\notin t)\rightarrow(s\in t\vee s=t\vee t\in s).$
If $s$ is a transitive set with transitive members then $\bigcup s$ is a transitive set with transitive members and $s\notin s\rightarrow\bigcup s\notin\bigcup s$. And $s\notin\bigcup s$ otherwise $s\in s$. Therefore $\bigcup s=s$ or $s=\bigcup s\cup\{\bigcup s\}$.
It is straightforward to define a number sequence satisfying the Peano Axioms for natural numbers. Natural numbers are also known as ‘finite ordinals’.
Definition 3.
Set $s$ is a ‘natural number with first number $\alpha$’ if $s$ is a transitive
set with transitive members, $\alpha$ is a pair of individuals and
$\bigcup\alpha=\alpha\wedge s\notin\alpha\wedge\forall u(u\in s\cup\{s\}\rightarrow(u\in\alpha\cup\{\alpha\}\vee\bigcup u\in u)).$
The ‘successor’ of natural number $s$ is $s\cup\{s\}$. $\Box$
If $z$ (‘zero’), $x$ and $y$ are individuals then there is number sequence $S$ with first number (in physical reality called a ‘unit of measurement’) $\{z,x\}$ (‘x coordinate’) and there is number sequence $T$ with first number (unit of measurement) $\{z,y\}$ (‘y coordinate’). All elements of both sequences have $z$ in common.
If $u$ belongs to the $S$ sequence or to the $T$ sequence then $x\in u$ implies $u$ is an $x$ coordinate and $y\in u$ implies $u$ is an $y$ coordinate.

### v4

@@
 Doeko H. Homan
-(Date: August 18, 2024)
-George Boolos describes the naive conception of set: ”Any predicate has an extension” (Boolos [1971] pp.215-231).
+(Date: January 15, 2025)
+George Boolos describes the naive conception of set ”Any predicate has an extension” (Boolos [1971] pp.215-231).
 Then Russell’s paradox is used to prove the inconsistency of naive set theory.
-Russell’s paradox reads:
+Russell’s paradox reads
 ”No set can contain all and only those sets which do not contain themselves”.
@@
 In (Boolos [1971] pp.216-217) ”individuals (= non-sets)”.
-In section 4.6.5 (Mendelson [2015]) Elliott Mendelson explains:
+In section 4.6.5 (Mendelson [2015]) Elliott Mendelson explains
 ”The words ‘individual’ and ‘atom’ are sometimes used as synonyms for
-‘urelement’. Thus, urelements have no members”. And in a footnote (p.304):
+‘urelement’. Thus, urelements have no members”. And in a footnote (p.304)
 ”Zermelo’s 1908 axiomatization permitted urelements”.
-In (Zermelo [1908] pp.262-265) Ernst Zermelo describes in ”10. Theorem.”:
+In (Zermelo [1908] pp.262-265) Ernst Zermelo describes in ”10. Theorem.”
 ”Für jedes Element $x$ von $M$ ist es definit, ob $x$ $\varepsilon$ $x$ oder nicht; diese Möglichkeit
@@
 Symbol $\in$ is the ‘membership relation’ symbol.
-Then $s\in t$ means $t$ contains $s$ or $s$ is a member of $t$.
+Then $s\in t$ means $s$ is ‘contained’ in $t$.
 The negation of $s\in t$ is denoted by $s\notin t$.
-The membership relation is definite, that is for every $s$ and for every $t$ ($s\in t$ or $s\notin t$) but never ($s\in t$ and $s\notin t$).
+The membership relation is definite, that is for every $s$ and for every $t$ either $s\in t$ or $s\notin t$ but not both.
 If $s\in s$ then $s$ is an individual.
-We prove Russell’s paradox.
+We prove Russell’s paradox
+”No set can contain all and only those sets which do not contain themselves”.
 Assume exists $s$ such that for every $u$ ($u\in s$ if and only if $u\notin u$).
@@
 The question ‘is $s$ contained in $s$’ or ‘is $s$ not contained in $s$’ is definite but there is
-no answer: If $s\in s$ then $s\notin s$ and vice versa. Thus $s$ is nonexisting. Therefore:
-”No set can contain all and only those sets which do not contain themselves”. $\Box$
-The membership relation is definite thus for every $s$ ($s\in s$ or $s\notin s$).
+no answer: If $s\in s$ then $s\notin s$ and vice versa. Therefore $s$ is nonexisting. $\Box$
+For every $s$ and for every $t$ the membership relation $s\in t$ is definite. Thus
 if $s\in s$ then exists $u$ such that ($u\in s$ and $u\in u$),
-if $s\notin s$ then exists $u$ such that ($u\notin s$ and $u\notin u$) therefore:
-For every $s$ exists $u$ such that (($u\in s$ and $u\in u$) or ($u\notin s$ and $u\notin u$)).
-That is logically equivalent to Russell’s paradox:
-It is not the case that exists $s$ such that for every $u$ ($u\in s$ if and only if $u\notin u$).
-Thus for every $s$, an individual is contained in $s$ or a set that is not an individual
-is not contained in $s$. $\Box$
+if $s\notin s$ then exists $u$ such that ($u\notin s$ and $u\notin u$). Then for every $s$, an individual is contained in $s$ or a set that is not an individual is not contained in $s$.
 Zermelo’s 1908 axiomatization permitted individuals.
@@
 But a set which contains itself, an individual, is problematic.
-(Boolos [1971] p.217, pp.219-220):
+(Boolos [1971] p.217, pp.219-220)
 ”perhaps the mind ought to boggle at the idea of something’s containing itself.”
@@
 ”Here are some things. Now we bind them up into a whole. Now we have a set.”
-However, ‘binding up some things’ or ‘containing something’ is not the definition of a set. (Boolos [1971] p.215):
+However, ‘binding up some things’ or ‘containing something’ is not the definition of a set. (Boolos [1971] p.215)
 ”Cantor also defined a set as a ‘many, which can be thought of as one, i.e. a
 totality of definite elements that can be combined into a whole by a law.’ ”
-Then ”which can be thought of as one,” a set is a ‘mental picture’ in your mind. Therefore we prefer to write ‘constituent of a set’ rather than ‘contained in a set’.
-Set theory is a language that follows the rules of logic to communicate the mental picture of a set with the physical world.
-Logic treats every possibility.
-Individuals are possible.
-Then, hopefully, someone reading or listening that language forms a correct mental picture in her/his mind.
-That is called ‘experience’.
-What is an individual? ‘Set $s$ is an individual’ is what is called in (Wittgenstein [1999]) ”an atomic fact”. In the Introduction (p.12) Bertrand Russell explains:
+Then ”which can be thought of as one,” thus a set is a ‘mental picture’ in your mind. Therefore we prefer to write ‘constituent of a set’, not ‘contained in a set’.
+Set theory is a language that follows rules of logic to communicate the mental picture of a set with the physical world.
+Logic treats every possibility, and $s\in s$ is possible.
+Then someone reading or listening that language forms a mental picture of a set in her/his mind.
+The picture in her/his mind agrees or disagrees with the picture in your mind.
+We call that ‘experience’.
+What is an individual? ‘Set $s$ is an individual’ is what is called in (Wittgenstein [1999]) ”an atomic fact”. In the Introduction (p.12) Bertrand Russell explains
 ”If an atomic fact is analyzed as fully as possible (theoretical, not practical
@@
 The members of a set are the constituents of that set.
-There is a membership relation between sets.
-Very basic set theory is based on
-the world of individuals,
-the membership relation is definite,
+Therefore set theory describes
+a world of individuals,
+with a definite membership relation between sets,
 the basic properties of a set are the constituents of that set and the sets that set is a member of,
@@
 Experience shows a set ‘is equal to’ or ‘is not equal to’ another set. Sets with different constituents are different sets, and different sets have different constituents.
-3. Very basic set theory
+Set theory follows rules of logic but the membership relation is definite. That means the logic of ‘true’ and ‘false’ should be changed to the logic of ‘definiteness’.
+In the next section we prove logical formulas preserve definiteness.
+3. Logic and definiteness
 A ‘formula’ is a description of properties of sets.
 Thus $s\in t$ is a formula.
-Formulas are denoted by $\Phi$, $\Psi$, $\ldots$ and are built up from formulas by means of negation $\neg\Phi$ (‘not $\Phi$’) and conjunction $\Phi\wedge\Psi$ (‘$\Phi$ and $\Psi$’).
-Negation has higher precedence than conjunction.
+Formulas are denoted by $\Phi$, $\Psi$, $\ldots$ and are built up from formulas by means of negation $\neg\Phi$ (not $\Phi$) and conjunction $\Phi\wedge\Psi$ ($\Phi$ and $\Psi$).
+Negation takes precedence over conjunction.
 The parenthesis (((( and )))) preclude ambiguity and improve legibility.
 Then $s\notin t$ is short for the formula $\neg(s\in t)$.
+If it is not the case that both $\Phi$ and $\neg\Phi$, that is formula $\neg(\Phi\wedge\neg\Phi)$, then ‘$\Phi$ is definite’.
+For every $s$ and for every $t$ it is not the case that both $s\in t$ and $\neg(s\in t)$ thus formula $s\in t$ is definite.
 If $\Phi$ is a formula then $(\Phi)$, $\forall s(\Phi)$ and $\exists s(\Phi)$ are formulas.
-$\forall s(\Phi)$ means ‘for every set $s$ $\Phi$’, and $\exists s(\Phi)$ means ‘exists set $s$ such that $\Phi$’.
-Then e.g.
-$\neg\exists s\forall u(\Phi)$ is short for the formula $\neg(\exists s(\forall u(\Phi)))$.
-If there is $s$ in $\Phi$ not guided in $\Phi$ by $\forall s$ or by $\exists s$ then ‘$s$ is free in $\Phi$’.
-If $s$ is free in $\Phi$ and $\neg\exists s(\Phi\wedge\neg\Phi)$ then ‘$\Phi$ is definite’.
-We define the logical connectives $\vee$, $\rightarrow$ and $\leftrightarrow$. The precedence of each of it is lower than conjunction.
-$\Phi\vee\Psi$ (‘$\Phi$ or $\Psi$ or both’) is short for the formula ¬($\neg(\neg\Phi\wedge\neg\Psi)$,
-$\Phi\rightarrow\Psi$ (‘$\Phi$ implies $\Psi$’) is short for the formula ($\neg\Phi\vee\Psi$),
-$\Phi\leftrightarrow\Psi$ (‘$\Phi$ if and only if $\Psi$’) is short for the formula ($(\Phi\rightarrow\Psi)\wedge(\Psi\rightarrow\Phi))$.
-If $\Phi$ and $\Psi$ are definite then $\neg\Phi$, $\Phi\wedge\Psi$, $\Phi\vee\Psi$, $\Phi\rightarrow\Psi$ and $\Phi\leftrightarrow\Psi$ are definite.
-Then Russell’s paradox is the formula $\neg\exists s\forall u(u\in s\leftrightarrow u\notin u)$, and is logically equivalent to $\forall s\exists u((u\in s\wedge u\in u)\vee(u\notin s\wedge u\notin u))$.
-The naive conception of set is inconsistent: formula $\exists s\forall u(u\in s\leftrightarrow u\notin u)$ is not definite. Therefore we postulate ‘specification’ (also known as ‘separation’).
-  • if $u$ is free in $\Phi$ and $\Phi$ is definite then $(u\in s\wedge\Phi)$ is definite and
-    (specification) $\forall s(\exists u(u\in s\wedge\Phi)\rightarrow\exists v\forall u(u\in v\leftrightarrow(u\in s\wedge\Phi)))$.
+$\forall s(\Phi)$ means ‘for every set it is the case $\Phi$’.
+The negation of $\forall s(\neg\Phi)$ is denoted by $\exists s(\Phi)$ thus $\exists s(\Phi)$ means ‘exists a set such that $\Phi$’.
+The parenthesis following $\forall s$ and $\exists s$ are mandatory.
+However, without ambiguity $\forall s\exists t(s\in t)$ is short for the formula $\forall s(\exists t(s\in t))$.
+Set $s$ in $\forall s$ or in $\exists s$ is not a ‘variable’.
+$\forall s$ means ‘for every set’, thus also for a set denoted by $s$.
+And $\exists s$ means ‘exists a set’ denoted by $s$.
+Set theory is a language.
+If there is $s$ in $\Phi$ not guided in $\Phi$ by $\forall s$ or by $\exists s$ then ‘$s$ is free in $\Phi$’. Thus $s$ is free and $t$ is not free in formula $\forall t(s\in t)$.
+If $\Phi$ is definite then $\forall s(\Phi)$ and $\exists s(\Phi)$ are definite otherwise exists $s$ such that $(\forall s(\Phi)\wedge\neg\forall s(\Phi))$ and $(\exists s(\Phi)\wedge\neg\exists s(\Phi))$ thus $(\Phi\wedge\neg\Phi)$.
+We define the logical connectives $\vee$, $\rightarrow$ and $\leftrightarrow$.
+$\Phi\vee\Psi$ ($\Phi$ or $\Psi$ or both) is short for the formula ¬($\neg(\neg\Phi\wedge\neg\Psi)$,
+$\Phi\rightarrow\Psi$ ($\Phi$ implies $\Psi$) is short for the formula ($(\neg\Phi\vee\Psi)$,
+$\Phi\leftrightarrow\Psi$ ($\Phi$ if and only if $\Psi$) is short for the formula ($(\Phi\rightarrow\Psi)\wedge(\Psi\rightarrow\Phi))$.
+There are many rewite rules: ¬($\neg(\neg\Phi)$ is logically equivalent to $\Phi$, $\neg(\Phi\leftrightarrow\Psi)$ is logically equivalent to $(\Phi\leftrightarrow\neg\Psi)$ and $((\Phi\wedge\Psi)\wedge\neg(\Phi\wedge\Psi))$ is logically equivalent to $((\Phi\wedge\Psi)\wedge(\neg\Phi\vee\neg\Psi))$.
+More rewrite rules in (Mendelson [2015] p.15).
+If $\Phi$ is definite then it is not the case that both $\Phi$ and $\neg\Phi$. Thus it is not the case that both ¬($\neg(\neg\Phi)$ and $\neg\Phi$ therefore $\neg\Phi$ is definite.
+And $(\Phi\wedge\Psi)\wedge\neg(\Phi\wedge\Psi)$ is logically equivalent to $((\Phi\wedge\neg\Phi)\wedge\Psi)\vee(\Phi\wedge(\Psi\wedge\neg\Psi))$ therefore if $\Phi$ and $\Psi$ are definite then $(\Phi\wedge\Psi)$ is definite.
+The logical ‘or’, ‘implies’ and ‘if and only if’ are expressed in terms of logical ‘not’ and logical ‘and’. Thus if $\Phi$ and $\Psi$ are definite then
+$\neg\Phi$, $\Phi\wedge\Psi$, $\Phi\vee\Psi$, $\Phi\rightarrow\Psi$, $\Phi\leftrightarrow\Psi$, $(\Phi)$, $\forall s(\Phi)$ and $\exists s(\Phi)$ are definite.
+Therefore the building up of formulas preserves definiteness.
+4. Russell’s paradox
+We prove formula $\neg\exists s\forall u(u\in s\leftrightarrow u\notin u)$, that is Russell’s paradox
+”No set can contain all and only those sets which do not contain themselves”.
+Assume exists set $s$ such that for every set $u$ $(u\in s\leftrightarrow u\notin u)$.
+The formula applies to every $u$ thus we can choose $s$ therefore
+exists $s$ such that $(s\in s\leftrightarrow s\notin s)$.
+Formula $(s\in s\leftrightarrow s\notin s)$ is short for $(s\in s\vee s\in s)\wedge(s\notin s\vee s\notin s)$ thus
+$(s\in s\wedge s\notin s)$.
+Formula $s\in s$ is definite thus it is not the case that $(s\in s\wedge s\notin s)$.
+Therefore set $s$ does not exists.
+Formula $\exists s\forall u(u\in s\leftrightarrow u\notin u)$ is definite therefore $\neg\exists s\forall u(u\in s\leftrightarrow u\notin u)$. $\Box$
+Then Russell’s paradox $\neg\exists s\forall u(u\in s\leftrightarrow u\notin u)$ is logically equivalent to
+$\forall s\exists u(u\in s\leftrightarrow u\in u)$ thus $\forall s\exists u((u\in s\wedge u\in u)\vee(u\notin u\wedge u\notin u))$.
+Therefore the logic of definiteness proves that
+for every set $s$, an individual is a constituent of $s$ or a set that is not an individual
+is not a constituent of $s$.
+5. Axiom ‘specification’
+If $\Phi$ is definite and if exists set $u\in s$ such that $\Phi$ then it is not the case that for every set $u$ $(u\in s\rightarrow\neg\Phi)$.
+Thus we can postulate a set $v$ whose constituents are all and only those constituents of $s$ such that for every constituent of $v$ it is the case $\Phi$.
+That is axiom specification (also known as ‘separation’).
+  • if $u$ is free in $\Phi$ and $\Phi$ is definite then
+    $\forall s(\exists u(u\in s\wedge\Phi)\rightarrow\exists v\forall u(u\in v\leftrightarrow(u\in s\wedge\Phi)))$.
 Thus the conjunction of mutually exclusive properties does not constitute a set.
-We prove: If $\exists u(u\in s\wedge u\in u)$ then exists $v$ whose members are all and only
-those individuals that are constituent of $s$.
-Formula $(u\in s\wedge u\in u)$ is definite and is not the conjunction of mutually exclusive properties. Apply ‘specification’ to $s$ to find $v$
+We prove: If an individual is a constituent of $s$ then exists $v$ whose constituents
+are all and only those individuals that are constituent of $s$.
+Formula $(u\in s\wedge u\in u)$ is definite and is not the conjunction of mutually exclusive properties. Apply ‘specification’ to set $s$ to find set $v$
 $\exists v\forall u(u\in v\leftrightarrow(u\in s\wedge u\in u))$. Then an individual is a constituent of $v$.
@@
 Then $\exists v(v\notin v\vee v\in s)$. Thus $v$ is not an individual, or $v$ is an individual and a constituent of $s$. A set of all individuals is not excluded. $\Box$
-We prove ‘the set of all sets’ is nonexisting.
-If $\exists u(u\in s\wedge u\notin u)$ then formula $(u\in s\wedge u\notin u)$ is definite and is not the conjuction of mutually exclusive properties.
-Apply ‘specification’ to $s$ to find $v$
-$\exists v\forall u(u\in v\leftrightarrow(u\in s\wedge u\notin u))$. Then a set is a member of $v$.
+We prove: The set of all sets is nonexisting.
+If $\exists u(u\in s\wedge u\notin u)$ then formula $(u\in s\wedge u\notin u)$ is definite and is not the conjuction of mutually exclusive properties. Apply ‘specification’ to set $s$ to find set $v$
+$\exists v\forall u(u\in v\leftrightarrow(u\in s\wedge u\notin u))$. Then a set is a constituent of $v$.
 The formula applies to every $u$ thus we can choose $v$
@@
 $\exists v((v\notin v\vee(v\in s\wedge v\notin v))\wedge((v\notin s\vee v\in v)\vee v\in v))$.
-Then $\exists v(v\notin v\wedge v\notin s)$. Thus $v$ is not an individual and $v$ is not a member of $s$. Therefore ‘the set of all sets’ is nonexisting. $\Box$
-4. More Axioms
-Experience shows a set is equal to or a set is not equal to another set.
-We define equality for sets.
+Then $\exists v(v\notin v\wedge v\notin s)$. Thus $v$ is not an individual and not a constituent of $s$. Therefore the set of all sets is nonexisting. $\Box$
+6. Axioms for ‘equality’
 Definition 1.
@@
 Then $\forall s\forall t((s\in s\wedge t\notin t)\rightarrow(s\neq t\wedge t\notin s))$.
+7. Axiom ‘pairs’
 Any pair of sets constitute a set. We postulate ‘pairs’.
@@
 $s\neq t\rightarrow\{s,t\}\notin\{s,t\}$. Thus a pair of different sets is not an individual.
-However, if $s\neq t$ and $(s\in t\wedge t\in s)$ then $s\notin s$ and $t\notin t$. Then exists $\{s,t\}$ and there is the ‘vicious membership circle’ $(s\in t\wedge t\in s)$.
+However, if $s\neq t$ and $(s\in t\wedge t\in s)$ then $s\notin s$ and $t\notin t$. Then exists set $\{s,t\}$ thus there is the ‘vicious membership circle’ $(s\in t\wedge t\in s)$.
+Therefore we postulate ‘regularity’.
+8. Axiom ‘regularity’
 The naming of individuals is logically first. Thus every set has a constituent, and vicious membership circles do not exist. Therefore we postulate ‘regularity’.
-  • (regularity) $\forall s\exists v(v\in s\wedge(v\notin v\rightarrow\forall u((u\in s\wedge u\in v)\rightarrow u\in u)))$.
+  • (regularity) $\forall s\exists v(v\in s\wedge\forall u((u\in s\wedge u\in v)\rightarrow u\in u))$.
 If $s$ is an individual or a set of individuals then $v$ is an individual. Otherwise exists $v$ and $v$ is not an individual and $s$ and $v$ only have individuals in common.
 And it is not the case that exists $s$ such that
-$\forall v(v\in s\wedge v\notin v\rightarrow\exists u(u\in s\wedge u\in v\wedge u\notin u))$.
-Therefore any sequence of sets $s$, $t$, $u$, $v$, … such that each term is a member of the previous one $t\in s$, $u\in t$, $v\in u$ is nonexisting.
-Thus $(t\in s\wedge s\in t)$ is nonexisting therefore $\forall s\forall t(s\neq t\rightarrow(s\notin t\vee t\notin s))$.
+$\forall v(v\in s\rightarrow\exists u((u\in s\wedge u\in v)\wedge u\notin u))$.
+Therefore any sequence of sets $s$, $t$, $u$, $v$, … , finite (a vicious membership circle) or not finite, such that each term is a member of the previous one
+$t\in s$, $u\in t$, $v\in u$ is nonexisting.
+Thus $\forall s\forall t(s\neq t\rightarrow(s\notin t\vee t\notin s))$.
+9. Axiom ‘union’
 The members of the members of a set constitute a set. We postulate ‘union’.
@@
 If $s$ and $t$ are individuals then $s\cup t=\{s,t\}$.
-An ‘infinite’ sequence of different sets $s$, $\bigcup s$, $\bigcup(\bigcup s)$, … is nonexisting.
-Any pair of different consecutive sets $\{w,\bigcup w\}$ implies $\exists v(v\in w\wedge v\notin v)$ otherwise $\bigcup w=w$. If $z=s\cup\bigcup s\cup\bigcup(\bigcup s)\ldots$ then
-$\forall w((w\in z\wedge w\notin w)\rightarrow\exists v(v\in w\wedge v\in\bigcup z\wedge v\notin v))$.
-Then $\forall v(v\in\bigcup z\rightarrow v\in z)$ thus $z$ contradicts regularity.
-Therefore the sequence terminates: $\exists w(\bigcup w=w)$. Set $z$ is called ‘the transitive closure’ of $s$. Thus for every set the transitive closure exists. $\Box$
-5. Conclusion
-Very basic set theory is a set theory with individuals
-the only constituent of an individual is that individual itself,
-Russell’s paradox is logically equivalent to a formula with individuals,
-the conjunction of mutually exclusive properties does not constitute a set,
-the naming of individuals is logically first,
+We prove: Any sequence of different sets $s$, $\bigcup s$, $\bigcup(\bigcup s)$, … is finite.
+Then $s\neq\bigcup s\rightarrow\exists t(t\in s\wedge t\notin t)$ otherwise $s=\bigcup s$.
+$\bigcup s\neq\bigcup(\bigcup s)\rightarrow\exists u\exists t(u\in t\wedge t\in s\wedge u\notin u\wedge t\notin t)$ otherwise $\bigcup s=\bigcup(\bigcup s)$,
+$\bigcup(\bigcup s)\neq\bigcup(\bigcup(\bigcup s))\rightarrow\exists v\exists u\exists t(v\in u\wedge u\in t\wedge t\in s\wedge v\notin v\wedge u\notin u\wedge t\notin t)$,
+Then there is an ‘infinite’ sequence of sets $s$, $t$, $u$, $v$, … such that each term is
+a member of the previous one $t\in s$, $u\in t$, $v\in u$, … contradictory to regularity.
+Therefore the sequence terminates thus exists $w$ belonging to the sequence and $\bigcup w=w$.
+If $z=s\cup\bigcup s\cup\bigcup(\bigcup s)\cup\ldots$ $w$ then $z$ is a set and $\bigcup z=z$.
+Set $z$ is called ‘the transitive closure’ of $s$.
 For every set the transitive closure exists.
-Compared with ‘mainstream set theory’ a big difference is ‘what is logically first’.
-The logic of set theory allows for individuals, and Russell’s paradox is logically equivalent to a formula with individuals.
-Therefore we choose for ‘the naming of individuals is logically first’, not for the postulate of an ‘empty set’.
-There are more differences. The naive conception of set is inconsistent thus axiom ‘specification’ is modified. The axiom ‘powerset’ is as usual but Cantor’s Theorem does not hold.
-  • (powerset) $\forall s\exists v\forall u(u\in v\leftrightarrow\forall t(t\in u\rightarrow t\in s))$.
-Set $v$ is denoted by ${\mathcal{P}}s$. Then ${\mathcal{P}}\{s\}=\{\{s\}\}$ thus the ‘number of members’ of $\{s\}$ is equal to the ‘number of members’ of ${\mathcal{P}}\{s\}$.
-Cantor’s Theorem (Zermelo [1908] p.276) reads: For every $s$,
-the ‘number of members’ of $s$ is less than the ‘number of members’ of ${\mathcal{P}}s$.
-Thus Cantor’s Theorem does not hold.
-The definition of equality and axiom ‘pairs’ are only meaningful in combination with sets that are logically first.
-Thus with axiom ‘regularity’.
-And ‘regularity’ is used with ‘union’ to prove that for every set the transitive closure exists.
-In the last section we define ‘number sequences’ satisfying the Peano Axioms (PA) for natural numbers.
-The first axiom of PA reads: There is a ‘first natural number’.
-Any pair of individuals is a first number of a natural number sequence.
-Therefore ‘the set of all natural numbers’ is nonexisting.
-However, a set of all natural numbers with a specified first number is not excluded.
-If in the world only one pair of individuals exists then indeed a set of all natural numbers with that pair of individuals as first number can postulated to exist.
-That is the way of ‘mainstream set theory’: There is no need for individuals, and the ‘empty set’ is postulated as ‘first natural number’.
+$\Box$
+The definition of equality and axiom pairs are only meaningful if exist sets the naming of which are logically first.
+Thus with individuals and axiom regularity.
+And regularity is used with axiom union to prove that for every set the transitive closure exists.
+The proof needs no other axioms and no ‘natural numbers’.
 10. Natural numbers
+In (Halmos [1974] p.47) the Peano Axioms for the set of all natural numbers ”used to be considered as the fountainhead of all mathematical knowledge”.
+In this section we define ‘number sequences’ satisfying the Peano Axioms for natural numbers.
 Definition 2.
@@
 Therefore $\exists v(v\in s\wedge v\notin v\wedge\forall u(u\in v\rightarrow u\in u))$. $\Box$
-If $u$ and $v$ are the only individuals belonging to transitive set $s$ then $\{u,v\}$ is the only possible set of individuals belonging to $s$. Therefore $\{u,v\}\in s\cup\{s\}$.
-If $s$ and $t$ are transitive sets with transitive members and the only individuals belonging to $s$ or to $t$ are $u$ and $v$, thus $\forall w((w\in s\cup t\wedge w\in w)\rightarrow w\in\{u,v\})$, then the ‘law of trichotomy’ for $s$ and $t$ is easily proved and reads
+If the only individuals that are constituent of transitive set $s$ are $u$ and $v$ then $\{u,v\}$ is the only possible set of individuals that is a member of $s$ therefore $\{u,v\}\in s\cup\{s\}$.
+If $s$ and $t$ are transitive sets with transitive members and the only individuals that are members of $s$ or of $t$ are $u$ and $v$, thus $\forall w((w\in s\cup t\wedge w\in w)\rightarrow w\in\{u,v\})$, then the ‘law of trichotomy’ for $s$ and $t$ is easily proved and reads
 $(s\notin s\wedge t\notin t)\rightarrow(s\in t\vee s=t\vee t\in s).$
-If $s$ is a transitive set with transitive members then $\bigcup s$ is a transitive set with transitive members and $s\notin s\rightarrow\bigcup s\notin\bigcup s$.
+If $s$ is a transitive set with transitive members then $\bigcup s$ is a transitive set with transitive members and $(s\notin s\rightarrow\bigcup s\notin\bigcup s)$.
 And $s\notin\bigcup s$ otherwise $s\in s$.
 Therefore $\bigcup s=s$ or $s=\bigcup s\cup\{\bigcup s\}$.
-It is straightforward to define a number sequence satisfying the Peano Axioms for natural numbers. Natural numbers are also known as ‘finite ordinals’.
+It is straightforward to define a number sequence satisfying the Peano Axioms (PA) for natural numbers.
+The first axiom of PA reads: There is a first natural number.
+Natural numbers are also known as ‘finite ordinals’.
 Definition 3.
@@
 The ‘successor’ of natural number $s$ is $s\cup\{s\}$. $\Box$
-If $z$ (‘zero’), $x$ and $y$ are individuals then there is number sequence $S$ with first number (in physical reality called a ‘unit of measurement’) $\{z,x\}$ (‘x coordinate’) and there is number sequence $T$ with first number (unit of measurement) $\{z,y\}$ (‘y coordinate’).
+Then $\alpha$ is a natural number.
+In physical reality a first number is called ‘the unit of measurement’.
+The classical Greek concept of number (arithmos) is ”a finite plurality composed of units, where a unit is whatever counts as one thing in the number under consideration.” (Mayberry [2000] p.21).
+If $z$ (‘zero’), $x$ and $y$ are individuals then there is number sequence $S$ with first number $\{z,x\}$ (‘x coordinate’) and there is number sequence $T$ with first number $\{z,y\}$ (‘y coordinate’).
 All elements of both sequences have $z$ in common.
-If $u$ belongs to the $S$ sequence or to the $T$ sequence then $x\in u$ implies $u$ is an $x$ coordinate and $y\in u$ implies $u$ is an $y$ coordinate.
+If $u$ belongs to the $S$ sequence then $x\in u$ implies $u$ is an $x$ coordinate, if $v$ belongs to the $T$ sequence then $y\in v$ implies $v$ is an $y$ coordinate.
+Then $\{u,v\}$ is an ‘ordered pair’ with an intrinsic intuitive meaning.
+The Kuratowski ordered pair does not have any intrinsic intuitive meaning.
+It is ”just a convenient way” to define ordered pairs (Mendelson [2015] p.235, Kanamori [2003] §5).
+Any pair of individuals is a first number of a natural number sequence.
+Therefore ‘the set of all natural numbers’ is nonexisting.
+However, to postulate ‘set $\omega$ of all natural numbers with first number $\alpha$’ is not excluded.
+Then $\bigcup\omega=\omega$.
+And if $\bigcup\beta=\beta$ we can define ‘transfinite ordinal numbers with first number $\beta$’ analogous to natural numbers (finite ordinal numbers) with first number $\alpha$.
+In the Introduction (Wittgenstein [1999] p.21) Bertrand Russell writes
+”No logic can be considered adequate until it has been shown to be capable of dealing with transfinite numbers”.
+Thus the logic in Very basic set theory is adequate.
+Conclusion
+Very basic set theory is a set theory with individuals, the logic of definiteness is adequate and the naming of individuals is logically first.
+But individuals are ‘not needed for mathematical purposes’ (Oliver and Smliley [2006] §5.1).
+Thus Russell’s paradox is not the result of the logic of definiteness otherwise individuals exist, and the naming of a special set is logically first (an ‘empty set’, or an ‘infinite set’).
+Individuals are banished from ‘mainstream set theory’.
+There are more differences with ‘mainstream set theory’. For example, the axiom ‘powerset’ is as usual but Cantor’s Theorem does not hold.
+  • (powerset) $\forall s\exists v\forall u(u\in v\leftrightarrow u\subseteq s)$.
+Set $v$ is denoted by ${\mathcal{P}}s$. Then ${\mathcal{P}}\{s\}=\{\{s\}\}$ thus the ‘number of members’ of $\{s\}$ is equal to the ‘number of members’ of ${\mathcal{P}}\{s\}$.
+Cantor’s Theorem (Zermelo [1908] p.276) reads: For every set $s$,
+the ‘number of members’ of $s$ is less than the ‘number of members’ of ${\mathcal{P}}s$.
+Thus Cantor’s Theorem does not hold.

### v5

@@
 Doeko H. Homan
-(Date: January 15, 2025)
+(Date: February 5, 2025)
 George Boolos describes the naive conception of set ”Any predicate has an extension” (Boolos [1971] pp.215-231).
@@
 is not contained in $s$.
-Therefore Very basic set theory is a set theory with individuals.
-1. What are individuals?
+Very basic set theory is a set theory with individuals that has its philosophical foundation in Ludwig Wittgenstein’s Tractatus Logico-Philosophicus.
+Introduction
 In (Kanamori [2003] pp.285-288) Akihiro Kanamori describes that with the ‘empty
@@
 Symbol $\in$ is the ‘membership relation’ symbol.
-Then $s\in t$ means $s$ is ‘contained’ in $t$.
+Then $s\in t$ means $s$ is a ‘constituent’ (also called a member) of set $t$.
 The negation of $s\in t$ is denoted by $s\notin t$.
-The membership relation is definite, that is for every $s$ and for every $t$ either $s\in t$ or $s\notin t$ but not both.
-If $s\in s$ then $s$ is an individual.
-We prove Russell’s paradox
+The membership relation is definite, that means
+  • for every $s$ and for every $t$ applies either $s\in t$ or $s\notin t$, but not both.
+Then for every $s$ it is the case $s\in s$ or $\notin s$. We prove Russell’s paradox
 ”No set can contain all and only those sets which do not contain themselves”.
@@
 no answer: If $s\in s$ then $s\notin s$ and vice versa. Therefore $s$ is nonexisting. $\Box$
-For every $s$ and for every $t$ the membership relation $s\in t$ is definite. Thus
-if $s\in s$ then exists $u$ such that ($u\in s$ and $u\in u$),
-if $s\notin s$ then exists $u$ such that ($u\notin s$ and $u\notin u$). Then for every $s$, an individual is contained in $s$ or a set that is not an individual is not contained in $s$.
-Zermelo’s 1908 axiomatization permitted individuals.
-And the prove of Russell’s paradox includes individuals and other sets.
-But a set which contains itself, an individual, is problematic.
-(Boolos [1971] p.217, pp.219-220)
+1. What is an individual?
+In set theory $s\in s$ is possible thus individuals are possible. But a set that contains itself, an individual, is problematic (Boolos [1971] p.217, pp.219-220)
 ”perhaps the mind ought to boggle at the idea of something’s containing itself.”
@@
 ”Here are some things. Now we bind them up into a whole. Now we have a set.”
-However, ‘binding up some things’ or ‘containing something’ is not the definition of a set. (Boolos [1971] p.215)
-”Cantor also defined a set as a ‘many, which can be thought of as one, i.e. a
-totality of definite elements that can be combined into a whole by a law.’ ”
-Then ”which can be thought of as one,” thus a set is a ‘mental picture’ in your mind. Therefore we prefer to write ‘constituent of a set’, not ‘contained in a set’.
+However, ‘binding up some things’ or ‘containing something’ is not the definition of a set. A set does not ‘contain’ something, the constituents of a set exist.
+Georg Cantor defined a set as (Boolos [1971] p.215)
+  • ”a ‘many, which can be thought of as one, i.e. a totality of definite
+    elements that can be combined into a whole by a law’.”
+Thus a set is something which can be thought of as one, therefore a set is a ‘mental picture’ in your mind. That is in line with (Wittgenstein [1999] p.39) proposition
+  • ”2.1 We make to ourselves pictures of facts.”
+Therefore we prefer to write ‘constituent of a set’, not ‘contained in a set’.
 Set theory is a language that follows rules of logic to communicate the mental picture of a set with the physical world.
-Logic treats every possibility, and $s\in s$ is possible.
-Then someone reading or listening that language forms a mental picture of a set in her/his mind.
-The picture in her/his mind agrees or disagrees with the picture in your mind.
-We call that ‘experience’.
-What is an individual? ‘Set $s$ is an individual’ is what is called in (Wittgenstein [1999]) ”an atomic fact”. In the Introduction (p.12) Bertrand Russell explains
+Then someone reading, seeing or listening that language forms a mental picture of a set.
+That picture agrees or disagrees with the picture in your mind.
+(Wittgenstein [1999] p.43) proposition
+  • ”2.222 In the agreement or disagreement of its sense with reality, its truth
+    or falsity consists.”
+An individual is what is called in (Wittgenstein [1999]) ”a simple sign” (propositions
+3.2, 3.201, 3.202, 3.26). In the Introduction (p.12) Bertrand Russell explains
 ”If an atomic fact is analyzed as fully as possible (theoretical, not practical
@@
 Experience shows a set ‘is equal to’ or ‘is not equal to’ another set. Sets with different constituents are different sets, and different sets have different constituents.
-Set theory follows rules of logic but the membership relation is definite. That means the logic of ‘true’ and ‘false’ should be changed to the logic of ‘definiteness’.
-In the next section we prove logical formulas preserve definiteness.
+Set theory follows rules of logic.
+Logic treats every possibility thus logic treats the definiteness of the membership relation.
+In the next section we prove the building up of logical formulas preserves definiteness.
 3. Logic and definiteness
@@
 If there is $s$ in $\Phi$ not guided in $\Phi$ by $\forall s$ or by $\exists s$ then ‘$s$ is free in $\Phi$’. Thus $s$ is free and $t$ is not free in formula $\forall t(s\in t)$.
-If $\Phi$ is definite then $\forall s(\Phi)$ and $\exists s(\Phi)$ are definite otherwise exists $s$ such that $(\forall s(\Phi)\wedge\neg\forall s(\Phi))$ and $(\exists s(\Phi)\wedge\neg\exists s(\Phi))$ thus $(\Phi\wedge\neg\Phi)$.
+If $\Phi$ is definite then $\forall s(\Phi)$ and $\exists s(\Phi)$ are definite otherwise $(\forall s(\Phi)\wedge\neg\forall s(\Phi))$ or $(\exists s(\Phi)\wedge\neg\exists s(\Phi))$ thus $\exists s(\Phi\wedge\neg\Phi)$.
 We define the logical connectives $\vee$, $\rightarrow$ and $\leftrightarrow$.
@@
 $\Phi\leftrightarrow\Psi$ ($\Phi$ if and only if $\Psi$) is short for the formula ($(\Phi\rightarrow\Psi)\wedge(\Psi\rightarrow\Phi))$.
-There are many rewite rules: ¬($\neg(\neg\Phi)$ is logically equivalent to $\Phi$, $\neg(\Phi\leftrightarrow\Psi)$ is logically equivalent to $(\Phi\leftrightarrow\neg\Psi)$ and $((\Phi\wedge\Psi)\wedge\neg(\Phi\wedge\Psi))$ is logically equivalent to $((\Phi\wedge\Psi)\wedge(\neg\Phi\vee\neg\Psi))$.
+There are many rewrite rules: ¬($\neg(\neg\Phi)$ is logically equivalent to $\Phi$, $\neg(\Phi\leftrightarrow\Psi)$ is logically equivalent to $(\Phi\leftrightarrow\neg\Psi)$ and $((\Phi\wedge\Psi)\wedge\neg(\Phi\wedge\Psi))$ is logically equivalent to $((\Phi\wedge\Psi)\wedge(\neg\Phi\vee\neg\Psi))$.
 More rewrite rules in (Mendelson [2015] p.15).
@@
 6. Axioms for ‘equality’
-Definition 1.
+Definition.
 Set $s$ ‘is equal to’ set $t$, denoted by $s=t$ if $\forall u(u\in s\leftrightarrow u\in t)$.
@@
 And regularity is used with axiom union to prove that for every set the transitive closure exists.
-The proof needs no other axioms and no ‘natural numbers’.
+The proof needs no other axioms and no ‘ set of all natural numbers’.
 10. Natural numbers
@@
 In this section we define ‘number sequences’ satisfying the Peano Axioms for natural numbers.
-Definition 2.
+Definition.
 Set $s$ is a ‘subset’ of $t$, denoted by $s\subseteq t$, if $\forall u(u\in s\rightarrow u\in t)$.
@@
 Every individual and every set of individuals is transitive. If $s$ is transitive then $\bigcup s$ and $s\cup\{s\}$ are transitive.
-Theorem 1.
+Theorem.
 If $s$ is transitive and $\exists u(u\in s\wedge u\notin u)$ then a set of individuals is a
@@
 Natural numbers are also known as ‘finite ordinals’.
-Definition 3.
+Definition.
 Set $s$ is a ‘natural number with first number $\alpha$’ if $s$ is a transitive
@@
 Conclusion
-Very basic set theory is a set theory with individuals, the logic of definiteness is adequate and the naming of individuals is logically first.
-But individuals are ‘not needed for mathematical purposes’ (Oliver and Smliley [2006] §5.1).
-Thus Russell’s paradox is not the result of the logic of definiteness otherwise individuals exist, and the naming of a special set is logically first (an ‘empty set’, or an ‘infinite set’).
-Individuals are banished from ‘mainstream set theory’.
-There are more differences with ‘mainstream set theory’. For example, the axiom ‘powerset’ is as usual but Cantor’s Theorem does not hold.
+Very basic set theory is a set theory that has its philosophical foundation in Ludwig Wittgenstein’s Tractatus Logico-Philosophicus.
+It is a set theory with individuals, the logic of definiteness is adequate and the naming of individuals is logically first.
+But individuals are not needed for mathematical purposes therefore individuals are banished from ‘mainstream set theory’.
+There are more differences with mainstream set theory. For example, the axiom ‘powerset’ is as usual but Cantor’s Theorem does not hold.
   • (powerset) $\forall s\exists v\forall u(u\in v\leftrightarrow u\subseteq s)$.

### v6

@@
 Very Basic Set theory
-Very Basic Set theory
-Doeko H. Homan
-(Date: February 5, 2025)
-George Boolos describes the naive conception of set ”Any predicate has an extension” (Boolos [1971] pp.215-231).
+Abstract.
+The naive conception of set is described as ”Any predicate has an extension”.
 Then Russell’s paradox is used to prove the inconsistency of naive set theory.
-Russell’s paradox reads
-”No set can contain all and only those sets which do not contain themselves”.
+Russell’s paradox reads ”No set can contain all and only those sets which do not contain themselves”.
 We call a set which contains itself an ‘individual’.
-In this article we show Russell’s paradox is logically equivalent to:
-For every set $s$, an individual is contained in $s$ or a set that is not an individual
-is not contained in $s$.
+In this article we show Russell’s paradox is logically equivalent to ‘for every set $s$, an individual is contained in $s$ or a set that is not an individual is not contained in $s$’.
 Very basic set theory is a set theory with individuals that has its philosophical foundation in Ludwig Wittgenstein’s Tractatus Logico-Philosophicus.
 Introduction
+In (Oliver and Smiley [2006] pp.145-146) Alex Oliver and Timothy Smiley explain
+individuals are not needed therefore are prohibited in set theory.
 In (Kanamori [2003] pp.285-288) Akihiro Kanamori describes that with the ‘empty
@@
 null set yet having no members and capable of belonging to sets”.
-In (Boolos [1971] pp.216-217) ”individuals (= non-sets)”.
+In (Boolos [1971] pp.216-217) George Boolos notes ”individuals (= non-sets)”.
 In section 4.6.5 (Mendelson [2015]) Elliott Mendelson explains
@@
 $x$ $\varepsilon$ $x$ ist an und für sich durch unsere Axiome nicht ausgeschlossen.”
-Thus ‘$x$ $\varepsilon$ $x$ or not’ is definite; the possibility ‘$x$ $\varepsilon$ $x$’ is not excluded by our axioms.
+Thus ‘$x$ $\varepsilon$ $x$ or not’ is definite; the possibility ”$x$ $\varepsilon$ $x$” is not excluded by our axioms.
 Therefore Zermelo’s 1908 axiomatization permitted sets which contain themselves.
@@
   • for every $s$ and for every $t$ applies either $s\in t$ or $s\notin t$, but not both.
-Then for every $s$ it is the case $s\in s$ or $\notin s$. We prove Russell’s paradox
-”No set can contain all and only those sets which do not contain themselves”.
-Assume exists $s$ such that for every $u$ ($u\in s$ if and only if $u\notin u$).
-The formula applies to every $u$ thus we can choose $s$ therefore
-exists $s$ such that ($s\in s$ if and only if $s\notin s$).
-The question ‘is $s$ contained in $s$’ or ‘is $s$ not contained in $s$’ is definite but there is
-no answer: If $s\in s$ then $s\notin s$ and vice versa. Therefore $s$ is nonexisting. $\Box$
+Then for every set $s$ it is the case $s\in s$ or $s\notin s$, but not both. We define
+  • a set is an ‘individual’ if the only constituent of that set is that set itself.
+In Section 1 and Section 2 we describe individuals and sets, in Section 3 we prove logical formulas preserve definiteness.
+Then Russell’s paradox is proved and axioms specification, equality, pairs, regularity and union are formulated.
+Some important theorems are proved, and in Section 10 natural numbers are defined.
 1. What is an individual?
@@
 The members of a set are the constituents of that set.
-Therefore set theory describes
+Therefore Very basic set theory describes
 a world of individuals,
@@
 Then $s\notin t$ is short for the formula $\neg(s\in t)$.
-If it is not the case that both $\Phi$ and $\neg\Phi$, that is formula $\neg(\Phi\wedge\neg\Phi)$, then ‘$\Phi$ is definite’.
-For every $s$ and for every $t$ it is not the case that both $s\in t$ and $\neg(s\in t)$ thus formula $s\in t$ is definite.
 If $\Phi$ is a formula then $(\Phi)$, $\forall s(\Phi)$ and $\exists s(\Phi)$ are formulas.
@@
 However, without ambiguity $\forall s\exists t(s\in t)$ is short for the formula $\forall s(\exists t(s\in t))$.
+If there is $s$ in $\Phi$ not guided in $\Phi$ by $\forall s$ or by $\exists s$ then ‘$s$ is free in $\Phi$’. Thus $s$ is free and $t$ is not free in formula $\forall t(s\in t)$.
 Set $s$ in $\forall s$ or in $\exists s$ is not a ‘variable’.
@@
 Set theory is a language.
-If there is $s$ in $\Phi$ not guided in $\Phi$ by $\forall s$ or by $\exists s$ then ‘$s$ is free in $\Phi$’. Thus $s$ is free and $t$ is not free in formula $\forall t(s\in t)$.
-If $\Phi$ is definite then $\forall s(\Phi)$ and $\exists s(\Phi)$ are definite otherwise $(\forall s(\Phi)\wedge\neg\forall s(\Phi))$ or $(\exists s(\Phi)\wedge\neg\exists s(\Phi))$ thus $\exists s(\Phi\wedge\neg\Phi)$.
 We define the logical connectives $\vee$, $\rightarrow$ and $\leftrightarrow$.
@@
 More rewrite rules in (Mendelson [2015] p.15).
-If $\Phi$ is definite then it is not the case that both $\Phi$ and $\neg\Phi$. Thus it is not the case that both ¬($\neg(\neg\Phi)$ and $\neg\Phi$ therefore $\neg\Phi$ is definite.
-And $(\Phi\wedge\Psi)\wedge\neg(\Phi\wedge\Psi)$ is logically equivalent to $((\Phi\wedge\neg\Phi)\wedge\Psi)\vee(\Phi\wedge(\Psi\wedge\neg\Psi))$ therefore if $\Phi$ and $\Psi$ are definite then $(\Phi\wedge\Psi)$ is definite.
-The logical ‘or’, ‘implies’ and ‘if and only if’ are expressed in terms of logical ‘not’ and logical ‘and’. Thus if $\Phi$ and $\Psi$ are definite then
-$\neg\Phi$, $\Phi\wedge\Psi$, $\Phi\vee\Psi$, $\Phi\rightarrow\Psi$, $\Phi\leftrightarrow\Psi$, $(\Phi)$, $\forall s(\Phi)$ and $\exists s(\Phi)$ are definite.
-Therefore the building up of formulas preserves definiteness.
+We define
+  • $\Phi$ is definite’ if $\neg(\Phi\wedge\neg\Phi)$ thus it is not the case that both $\Phi$ and $\neg\Phi$.
+For every $s$ and for every $t$ it is not the case that both $s\in t$ and $\neg(s\in t)$ thus $s\in t$ is definite We prove: The building up of logical formulas preserves definiteness.
+If $\Phi$ is definite then $\forall s(\Phi)$ and $\exists s(\Phi)$ are definite otherwise
+$(\forall s(\Phi)\wedge\neg\forall s(\Phi))$ or $(\exists s(\Phi)\wedge\neg\exists s(\Phi))$ thus $\exists s(\Phi\wedge\neg\Phi)$,
+If $\Phi$ is definite then it is not the case that both $\Phi$ and $\neg\Phi$ thus
+it is not the case that both ¬($\neg(\neg\Phi)$ and $\neg\Phi$ therefore $\neg\Phi$ is definite,
+$(\Phi\wedge\Psi)\wedge\neg(\Phi\wedge\Psi)$ is logically equivalent to $((\Phi\wedge\neg\Phi)\wedge\Psi)\vee(\Phi\wedge(\Psi\wedge\neg\Psi))$ thus if $\Phi$ and $\Psi$ are definite then $(\Phi\wedge\Psi)$ is definite,
+The connectives $\vee$, $\rightarrow$ and $\leftrightarrow$ are expressed in terms of $\neg$ and $\wedge$ thus if $\Phi$ and $\Psi$ are definite then $\Phi\vee\Psi$, $\Phi\rightarrow\Psi$ and $\Phi\leftrightarrow\Psi$ are definite.
+Therefore the building up of logical formulas preserves definiteness. $\Box$
 4. Russell’s paradox
-We prove formula $\neg\exists s\forall u(u\in s\leftrightarrow u\notin u)$, that is Russell’s paradox
+We prove Russell’s paradox, that is formula $\neg\exists s\forall u(u\in s\leftrightarrow u\notin u)$
 ”No set can contain all and only those sets which do not contain themselves”.
@@
 Then Russell’s paradox $\neg\exists s\forall u(u\in s\leftrightarrow u\notin u)$ is logically equivalent to
-$\forall s\exists u(u\in s\leftrightarrow u\in u)$ thus $\forall s\exists u((u\in s\wedge u\in u)\vee(u\notin u\wedge u\notin u))$.
+$\forall s\exists u(u\in s\leftrightarrow u\in u)$ thus $\forall s\exists u((u\in s\wedge u\in u)\vee(u\notin s\wedge u\notin u))$.
 Therefore the logic of definiteness proves that
@@
 $s\neq t\rightarrow\{s,t\}\notin\{s,t\}$. Thus a pair of different sets is not an individual.
-However, if $s\neq t$ and $(s\in t\wedge t\in s)$ then $s\notin s$ and $t\notin t$. Then exists set $\{s,t\}$ thus there is the ‘vicious membership circle’ $(s\in t\wedge t\in s)$.
-Therefore we postulate ‘regularity’.
+However, if $s\neq t$ and $(s\in t\wedge t\in s)$ then $s\notin s$ and $t\notin t$.
+Then exists set $\{s,t\}$ thus there is the ‘vicious membership circle’ $(s\in t\wedge t\in s)$.
+That is contradictory to ‘the naming of individuals is logically first’.
 8. Axiom ‘regularity’
@@
 If $s$ is an individual or a set of individuals then $v$ is an individual. Otherwise exists $v$ and $v$ is not an individual and $s$ and $v$ only have individuals in common.
-And it is not the case that exists $s$ such that
+Then it is not the case that exists $s$ such that
 $\forall v(v\in s\rightarrow\exists u((u\in s\wedge u\in v)\wedge u\notin u))$.
 Therefore any sequence of sets $s$, $t$, $u$, $v$, … , finite (a vicious membership circle) or not finite, such that each term is a member of the previous one
-$t\in s$, $u\in t$, $v\in u$ is nonexisting.
-Thus $\forall s\forall t(s\neq t\rightarrow(s\notin t\vee t\notin s))$.
+$t\in s$, $u\in t$, $v\in u$ is contradictory to ‘the naming of individuals is logically first’. Then $\forall s\forall t(s\neq t\rightarrow(s\notin t\vee t\notin s))$.
 9. Axiom ‘union’
@@
 Therefore $\bigcup s=s$ or $s=\bigcup s\cup\{\bigcup s\}$.
-It is straightforward to define a number sequence satisfying the Peano Axioms (PA) for natural numbers.
-The first axiom of PA reads: There is a first natural number.
-Natural numbers are also known as ‘finite ordinals’.
+It is straightforward to define a number sequence satisfying the Peano Axioms for natural numbers.
+The first axiom of the Peano Axioms reads ‘there is a first natural number’ (denoted in (Mendelson [2015] p.153) by ‘00’).
 Definition.
@@
 The ‘successor’ of natural number $s$ is $s\cup\{s\}$. $\Box$
-Then $\alpha$ is a natural number.
+Then $\alpha$ is a first natural number.
 In physical reality a first number is called ‘the unit of measurement’.
@@
 Then $\bigcup\omega=\omega$.
-And if $\bigcup\beta=\beta$ we can define ‘transfinite ordinal numbers with first number $\beta$’ analogous to natural numbers (finite ordinal numbers) with first number $\alpha$.
+Therefore we can define ‘transfinite ordinal numbers with first number $\omega$’ analogous to natural numbers (the finite ordinal numbers) with first number $\alpha$.
 In the Introduction (Wittgenstein [1999] p.21) Bertrand Russell writes
 ”No logic can be considered adequate until it has been shown to be capable of dealing with transfinite numbers”.
-Thus the logic in Very basic set theory is adequate.
+The logic in Very basic set theory is capable to deal with transfinite numbers.
 Conclusion
 Very basic set theory is a set theory that has its philosophical foundation in Ludwig Wittgenstein’s Tractatus Logico-Philosophicus.
-It is a set theory with individuals, the logic of definiteness is adequate and the naming of individuals is logically first.
+It is a set theory with individuals, the logic is capable to deal with transfinite numbers and the naming of individuals is logically first.
 But individuals are not needed for mathematical purposes therefore individuals are banished from ‘mainstream set theory’.
@@
   • (powerset) $\forall s\exists v\forall u(u\in v\leftrightarrow u\subseteq s)$.
-Set $v$ is denoted by ${\mathcal{P}}s$. Then ${\mathcal{P}}\{s\}=\{\{s\}\}$ thus the ‘number of members’ of $\{s\}$ is equal to the ‘number of members’ of ${\mathcal{P}}\{s\}$.
-Cantor’s Theorem (Zermelo [1908] p.276) reads: For every set $s$,
-the ‘number of members’ of $s$ is less than the ‘number of members’ of ${\mathcal{P}}s$.
-Thus Cantor’s Theorem does not hold.
+Set $v$ is denoted by ${\mathcal{P}}s$. Then ${\mathcal{P}}\{s\}=\{\{s\}\}$ thus the ‘number of members’ of $\{s\}$
+is equal to the ‘number of members’ of ${\mathcal{P}}\{s\}$.
+Cantor’s Theorem (Zermelo [1908] p.276) reads
+for every set $s$, the ‘number of members’ of $s$ is less than the ‘number of
+members’ of ${\mathcal{P}}s$.
+Thus Cantor’s Theorem does not hold. $\Box$

### v7

@@
 Then Russell’s paradox is used to prove the inconsistency of naive set theory.
-Russell’s paradox reads ”No set can contain all and only those sets which do not contain themselves”.
-We call a set which contains itself an ‘individual’.
-In this article we show Russell’s paradox is logically equivalent to ‘for every set $s$, an individual is contained in $s$ or a set that is not an individual is not contained in $s$’.
-Very basic set theory is a set theory with individuals that has its philosophical foundation in Ludwig Wittgenstein’s Tractatus Logico-Philosophicus.
+If the only constituent of a set is that set itself we call that set an ‘individual’.
+We show Russell’s paradox permits sets that are individuals thus that is not an inconsistency.
+Zermelo’s axiomatization also permits individuals.
+Very basic set theory is a set theory with individuals and has a philosophical foundation in Ludwig Wittgenstein’s Tractatus Logico-Philosophicus.
 Introduction
-In (Oliver and Smiley [2006] pp.145-146) Alex Oliver and Timothy Smiley explain
-individuals are not needed therefore are prohibited in set theory.
+Georg Cantor defined a set as (Boolos [1971] p.215)
+  • ”a ‘many, which can be thought of as one, i.e. a totality of definite
+    elements that can be combined into a whole by a law’.”
+That means a set is ‘determined’ by its elements.
+Sets are denoted by $s$, $t$, $u$, ….
+Symbol $\in$ is the ‘membership relation’ symbol.
+Then $s\in t$ means $s$ is a ‘constituent’ (also called ‘an element’ or ‘a member’) of the totality $t$.
+The negation of $s\in t$ is denoted by $s\notin t$.
+Set $s$ is an ‘individual’ if the only constituent of $s$ is $s$ itself.
+A set with only $s$ and $t$ as constituents is denoted by $\{s,t\}$ (Zermelo [1908] p.263).
+Often the terminology is: ‘set $\{s,t\}$ contains only $s$ and $t$’.
+But the totality $\{s,t\}$ is ‘thought of as one’ thus a set is a mental picture in your mind.
+Therefore we prefer to write ‘constituent of a set’, not ‘is contained in a set’.
+In (Oliver and Smiley [2006] $\S$5.1) Alex Oliver and Timothy Smiley explain why
+”numerous textbooks” prohibit individuals: individuals are not needed for
+mathematical purposes.
 In (Kanamori [2003] pp.285-288) Akihiro Kanamori describes that with the ‘empty
@@
 $x$ $\varepsilon$ $x$ ist an und für sich durch unsere Axiome nicht ausgeschlossen.”
-Thus ‘$x$ $\varepsilon$ $x$ or not’ is definite; the possibility ”$x$ $\varepsilon$ $x$” is not excluded by our axioms.
-Therefore Zermelo’s 1908 axiomatization permitted sets which contain themselves.
-Sets are denoted by $s$, $t$, $u$, ….
-Symbol $\in$ is the ‘membership relation’ symbol.
-Then $s\in t$ means $s$ is a ‘constituent’ (also called a member) of set $t$.
-The negation of $s\in t$ is denoted by $s\notin t$.
+Thus ‘$x$ $\varepsilon$ $x$ or not’ is definite; the possibility ”$x$ $\varepsilon$ xx” is not excluded by our axioms.
+Therefore Zermelo’s 1908 axiomatization permits sets ‘containing’ themselves.
 The membership relation is definite, that means
   • for every $s$ and for every $t$ applies either $s\in t$ or $s\notin t$, but not both.
-Then for every set $s$ it is the case $s\in s$ or $s\notin s$, but not both. We define
-  • a set is an ‘individual’ if the only constituent of that set is that set itself.
-In Section 1 and Section 2 we describe individuals and sets, in Section 3 we prove logical formulas preserve definiteness.
-Then Russell’s paradox is proved and axioms specification, equality, pairs, regularity and union are formulated.
-Some important theorems are proved, and in Section 10 natural numbers are defined.
-1. What is an individual?
-In set theory $s\in s$ is possible thus individuals are possible. But a set that contains itself, an individual, is problematic (Boolos [1971] p.217, pp.219-220)
+Then for every set $s$ it is the case $s\in s$ or $s\notin s$, but not both.
+Therefore $s\in s$ is possible.
+But a set that ‘contains’ itself is considered problematic (Boolos [1971] p.217, pp.219-220)
 ”perhaps the mind ought to boggle at the idea of something’s containing itself.”
@@
 ”Here are some things. Now we bind them up into a whole. Now we have a set.”
-However, ‘binding up some things’ or ‘containing something’ is not the definition of a set. A set does not ‘contain’ something, the constituents of a set exist.
-Georg Cantor defined a set as (Boolos [1971] p.215)
-  • ”a ‘many, which can be thought of as one, i.e. a totality of definite
-    elements that can be combined into a whole by a law’.”
-Thus a set is something which can be thought of as one, therefore a set is a ‘mental picture’ in your mind. That is in line with (Wittgenstein [1999] p.39) proposition
-  • ”2.1 We make to ourselves pictures of facts.”
-Therefore we prefer to write ‘constituent of a set’, not ‘contained in a set’.
-Set theory is a language that follows rules of logic to communicate the mental picture of a set with the physical world.
-Then someone reading, seeing or listening that language forms a mental picture of a set.
-That picture agrees or disagrees with the picture in your mind.
-(Wittgenstein [1999] p.43) proposition
-  • ”2.222 In the agreement or disagreement of its sense with reality, its truth
-    or falsity consists.”
-An individual is what is called in (Wittgenstein [1999]) ”a simple sign” (propositions
-3.2, 3.201, 3.202, 3.26). In the Introduction (p.12) Bertrand Russell explains
-”If an atomic fact is analyzed as fully as possible (theoretical, not practical
-possibility is meant) the constituents finally reached may be called simples
-or objects.”
-”(…) the naming of simples is shown to be what is logically first in logic.”
-Therefore in Very basic set theory the naming of individuals is logically first.
-2. What is a set?
+However, ‘binding up some things’ or ‘containing something’ is not the definition of a set. A set does not ‘contain’ something, the constituents of a set can be thought of as one totality.
+Russell’s paradox reads (Boolos [1971] p.217)
+”No set can contain all and only those sets which do not contain themselves.”
+The negation of Russell’s paradox reads
+  • for every $s$ exists a set $u$ such that (($u\in s$ and $u\in u$) or ($u\notin s$ and $u\notin u$)).
+If we choose for set $u$ set $s$ then the formula reads ‘for every $s$ ($s\in s$ or $s\notin s$)’.
+It is the case that ‘for every $s$ ($s\in s$)’ is definite.
+Therefore Russell’s paradox permits sets ‘containing’ themselves.
+In $\S$1 we describe sets and individuals.
+Atomic facts and elementary propositions are defined which gives the foundation of set theory in the Tractatus (Wittgenstein [1999]).
+In $\S$2 we prove the formulas of set theory preserve definiteness.
+In $\S$3 we define equality for sets.
+Then axioms specification, pairs and union are formulated.
+We prove regularity and some other important theorems.
+In $\S$7 natural numbers are defined.
+1. What are individuals and sets?
 At an early age you develop the mental picture of ‘what is a set’.
@@
 The members of a set are the constituents of that set.
-Therefore Very basic set theory describes
-a world of individuals,
-with a definite membership relation between sets,
-the basic properties of a set are the constituents of that set and the sets that set is a member of,
-the only constituent of an individual is that individual itself.
-Experience shows a set ‘is equal to’ or ‘is not equal to’ another set. Sets with different constituents are different sets, and different sets have different constituents.
-Set theory follows rules of logic.
-Logic treats every possibility thus logic treats the definiteness of the membership relation.
-In the next section we prove the building up of logical formulas preserves definiteness.
+Set theory is a language that follows rules of logic to communicate the mental picture of a set with the physical world.
+Then someone reading, seeing or listening that language forms a mental picture of a set.
+Set theory is made up of ‘formulas’.
+A formula is a proposition or predicate about the facts of sets.
+For every set $s$ and for every set $t$ the fact is the definiteness of $s\in t$.
+A formula which is not complex or not a combination of other formulas is an ‘atomic fact’.
+Formula $s\in t$ is an atomic fact, formula $t\in s$ is an atomic fact.
+Formula $s\notin t$ is the negation of $s\in t$ thus $s\notin t$ is a complex formula therefore $s\notin t$ is not an atomic fact.
+Atomic facts are what makes formulas true or false.
+Definition.
+Set $s$ is an individual if the only constituent of set $s$ is set $s$ itself.
+A formula presupposes propositions or predicates while propositions or predicates presupposes atomic facts.
+Then if an atomic fact is analyzed as fully as possible the constituents finally reached are individuals.
+Individuals cannot be analyzed further.
+In this way the naming of individuals is what is logically first in logic (Wittgenstein [1999] prop.
+3.26) and Bertrand Russell (Introduction p.12).
+If $s$ is a set then $s$ is not a proposition or predicate.
+If $s$ is an individual then $s\in s$ therefore $s$ is the ’propositional connexion’ (Wittgenstein [1999] prop.
+4.221), (Anscombe [1963] p.29-31).
+Thus individuals are ‘elementary propositions’.
+Therefore in Very basic set theory the naming of individuals is logically first.
 2. Logic and definiteness
-A ‘formula’ is a description of properties of sets.
-Thus $s\in t$ is a formula.
+In this section we prove logical formulas preserve definiteness.
 Formulas are denoted by $\Phi$, $\Psi$, $\ldots$ and are built up from formulas by means of negation $\neg\Phi$ (not $\Phi$) and conjunction $\Phi\wedge\Psi$ ($\Phi$ and $\Psi$).
 Negation takes precedence over conjunction.
-The parenthesis (((( and )))) preclude ambiguity and improve legibility.
-Then $s\notin t$ is short for the formula $\neg(s\in t)$.
-If $\Phi$ is a formula then $(\Phi)$, $\forall s(\Phi)$ and $\exists s(\Phi)$ are formulas.
+If $\Phi$ is a formula then $(\Phi)$ is a formula.
+The parenthesis (( and )) preclude ambiguity and improve legibility.
+Then $s\notin t$ is the formula $\neg(s\in t)$.
+If $\Phi$ is a formula then $\forall s(\Phi)$ and $\exists s(\Phi)$ are formulas.
 $\forall s(\Phi)$ means ‘for every set it is the case $\Phi$’.
@@
 The parenthesis following $\forall s$ and $\exists s$ are mandatory.
-However, without ambiguity $\forall s\exists t(s\in t)$ is short for the formula $\forall s(\exists t(s\in t))$.
+However, e.g.
+without ambiguity $\forall s\exists t(s\in t)$ is short for the formula $\forall s(\exists t(s\in t))$.
 If there is $s$ in $\Phi$ not guided in $\Phi$ by $\forall s$ or by $\exists s$ then ‘$s$ is free in $\Phi$’. Thus $s$ is free and $t$ is not free in formula $\forall t(s\in t)$.
@@
 We define the logical connectives $\vee$, $\rightarrow$ and $\leftrightarrow$.
-$\Phi\vee\Psi$ ($\Phi$ or $\Psi$ or both) is short for the formula ¬($\neg(\neg\Phi\wedge\neg\Psi)$,
-$\Phi\rightarrow\Psi$ ($\Phi$ implies $\Psi$) is short for the formula ($(\neg\Phi\vee\Psi)$,
+$\Phi\vee\Psi$ ($\Phi$ or $\Psi$ or both) is short for the formula $\neg(\neg\Phi\wedge\neg\Psi)$,
+$\Phi\rightarrow\Psi$ ($\Phi$ implies $\Psi$) is short for the formula $(\neg\Phi\vee\Psi)$,
 $\Phi\leftrightarrow\Psi$ ($\Phi$ if and only if $\Psi$) is short for the formula ($(\Phi\rightarrow\Psi)\wedge(\Psi\rightarrow\Phi))$.
-There are many rewrite rules: ¬($\neg(\neg\Phi)$ is logically equivalent to $\Phi$, $\neg(\Phi\leftrightarrow\Psi)$ is logically equivalent to $(\Phi\leftrightarrow\neg\Psi)$ and $((\Phi\wedge\Psi)\wedge\neg(\Phi\wedge\Psi))$ is logically equivalent to $((\Phi\wedge\Psi)\wedge(\neg\Phi\vee\neg\Psi))$.
+There are many rewrite rules: $\neg(\neg\Phi)$ is logically equivalent to $\Phi$, $\neg(\Phi\leftrightarrow\Psi)$ is logically equivalent to $(\Phi\leftrightarrow\neg\Psi)$ and $((\Phi\wedge\Psi)\wedge\neg(\Phi\wedge\Psi))$ is logically equivalent to $((\Phi\wedge\Psi)\wedge(\neg\Phi\vee\neg\Psi))$.
 More rewrite rules in (Mendelson [2015] p.15).
-We define
-  • $\Phi$ is definite’ if $\neg(\Phi\wedge\neg\Phi)$ thus it is not the case that both $\Phi$ and $\neg\Phi$.
-For every $s$ and for every $t$ it is not the case that both $s\in t$ and $\neg(s\in t)$ thus $s\in t$ is definite We prove: The building up of logical formulas preserves definiteness.
-If $\Phi$ is definite then $\forall s(\Phi)$ and $\exists s(\Phi)$ are definite otherwise
-$(\forall s(\Phi)\wedge\neg\forall s(\Phi))$ or $(\exists s(\Phi)\wedge\neg\exists s(\Phi))$ thus $\exists s(\Phi\wedge\neg\Phi)$,
-If $\Phi$ is definite then it is not the case that both $\Phi$ and $\neg\Phi$ thus
-it is not the case that both ¬($\neg(\neg\Phi)$ and $\neg\Phi$ therefore $\neg\Phi$ is definite,
-$(\Phi\wedge\Psi)\wedge\neg(\Phi\wedge\Psi)$ is logically equivalent to $((\Phi\wedge\neg\Phi)\wedge\Psi)\vee(\Phi\wedge(\Psi\wedge\neg\Psi))$ thus if $\Phi$ and $\Psi$ are definite then $(\Phi\wedge\Psi)$ is definite,
-The connectives $\vee$, $\rightarrow$ and $\leftrightarrow$ are expressed in terms of $\neg$ and $\wedge$ thus if $\Phi$ and $\Psi$ are definite then $\Phi\vee\Psi$, $\Phi\rightarrow\Psi$ and $\Phi\leftrightarrow\Psi$ are definite.
-Therefore the building up of logical formulas preserves definiteness. $\Box$
-4. Russell’s paradox
-We prove Russell’s paradox, that is formula $\neg\exists s\forall u(u\in s\leftrightarrow u\notin u)$
-”No set can contain all and only those sets which do not contain themselves”.
-Assume exists set $s$ such that for every set $u$ $(u\in s\leftrightarrow u\notin u)$.
-The formula applies to every $u$ thus we can choose $s$ therefore
-exists $s$ such that $(s\in s\leftrightarrow s\notin s)$.
-Formula $(s\in s\leftrightarrow s\notin s)$ is short for $(s\in s\vee s\in s)\wedge(s\notin s\vee s\notin s)$ thus
-$(s\in s\wedge s\notin s)$.
-Formula $s\in s$ is definite thus it is not the case that $(s\in s\wedge s\notin s)$.
-Therefore set $s$ does not exists.
-Formula $\exists s\forall u(u\in s\leftrightarrow u\notin u)$ is definite therefore $\neg\exists s\forall u(u\in s\leftrightarrow u\notin u)$. $\Box$
-Then Russell’s paradox $\neg\exists s\forall u(u\in s\leftrightarrow u\notin u)$ is logically equivalent to
-$\forall s\exists u(u\in s\leftrightarrow u\in u)$ thus $\forall s\exists u((u\in s\wedge u\in u)\vee(u\notin s\wedge u\notin u))$.
-Therefore the logic of definiteness proves that
-for every set $s$, an individual is a constituent of $s$ or a set that is not an individual
-is not a constituent of $s$.
-5. Axiom ‘specification’
-If $\Phi$ is definite and if exists set $u\in s$ such that $\Phi$ then it is not the case that for every set $u$ $(u\in s\rightarrow\neg\Phi)$.
-Thus we can postulate a set $v$ whose constituents are all and only those constituents of $s$ such that for every constituent of $v$ it is the case $\Phi$.
-That is axiom specification (also known as ‘separation’).
-  • if $u$ is free in $\Phi$ and $\Phi$ is definite then
-    $\forall s(\exists u(u\in s\wedge\Phi)\rightarrow\exists v\forall u(u\in v\leftrightarrow(u\in s\wedge\Phi)))$.
-Thus the conjunction of mutually exclusive properties does not constitute a set.
-We prove: If an individual is a constituent of $s$ then exists $v$ whose constituents
-are all and only those individuals that are constituent of $s$.
-Formula $(u\in s\wedge u\in u)$ is definite and is not the conjunction of mutually exclusive properties. Apply ‘specification’ to set $s$ to find set $v$
-$\exists v\forall u(u\in v\leftrightarrow(u\in s\wedge u\in u))$. Then an individual is a constituent of $v$.
-The formula applies to every $u$ thus we can choose $v$
-$\exists v(v\in v\leftrightarrow(v\in s\wedge v\in v))$ therefore
-$\exists v((v\notin v\vee(v\in s\wedge v\in v))\wedge((v\notin s\vee v\notin v)\vee v\in v))$.
-Then $\exists v(v\notin v\vee v\in s)$. Thus $v$ is not an individual, or $v$ is an individual and a constituent of $s$. A set of all individuals is not excluded. $\Box$
-We prove: The set of all sets is nonexisting.
-If $\exists u(u\in s\wedge u\notin u)$ then formula $(u\in s\wedge u\notin u)$ is definite and is not the conjuction of mutually exclusive properties. Apply ‘specification’ to set $s$ to find set $v$
-$\exists v\forall u(u\in v\leftrightarrow(u\in s\wedge u\notin u))$. Then a set is a constituent of $v$.
-The formula applies to every $u$ thus we can choose $v$
-$\exists v(v\in v\leftrightarrow(v\in s\wedge v\notin v))$ therefore
-$\exists v((v\notin v\vee(v\in s\wedge v\notin v))\wedge((v\notin s\vee v\in v)\vee v\in v))$.
-Then $\exists v(v\notin v\wedge v\notin s)$. Thus $v$ is not an individual and not a constituent of $s$. Therefore the set of all sets is nonexisting. $\Box$
-6. Axioms for ‘equality’
+Definition.
+Formula $\Phi$ is ‘definite’ if it is not the case that both $\Phi$ and $\neg\Phi$. $\Box$
+Theorem.
+The building up of formulas preserve definiteness.
+Proof. For every $s$ and for every $t$ it is not the case that both $s\in t$ and $\neg(s\in t)$
+therefore atomic facts and elementary propositions are definite.
+If $\Phi$ is definite then it is not the case that both $\Phi$ and $\neg\Phi$. Thus it is not the
+case that both $\neg(\neg\Phi)$ and $\neg\Phi$ therefore $\neg\Phi$ is definite.
+If $\Phi$ is definite then $\forall s(\Phi)$ and $\exists s(\Phi)$ are definite otherwise $(\forall s(\Phi)\wedge\neg\forall s(\Phi))$ or
+$(\exists s(\Phi)\wedge\neg\exists s(\Phi))$ thus $\exists s(\Phi\wedge\neg\Phi)$.
+And $(\Phi\wedge\Psi)\wedge\neg(\Phi\wedge\Psi)$ is logically equivalent to
+$((\Phi\wedge\neg\Phi)\wedge\Psi)\vee(\Phi\wedge(\Psi\wedge\neg\Psi))$ therefore if $\Phi$ and $\Psi$ are definite then
+$(\Phi\wedge\Psi)$ is definite.
+The logical connectives ‘or’, ‘implies’ and ‘if and only if’ are expressed in terms
+of logical ‘not’ and logical ‘and’. Thus if $\Phi$ and $\Psi$ are definite then $(\Phi\vee\Psi)$,
+$(\Phi\rightarrow\Psi)$ and $(\Phi\leftrightarrow\Psi)$ are definite.
+If an atomic fact is analyzed as fully as possible constituents finally reached are
+individuals. Atomic facts and elementary propositions are definite.
+Therefore the building up of formulas preserve definiteness. $\Box$
+3. Equality for sets
+Experience shows a set ‘is equal to’ or ‘is not equal to’ another set.
+A set is determined by its constituents.
+Thus sets with different constituents are different sets, and the constituents of different sets are different.
+We define equality for sets.
 Definition.
@@
 Then $\forall s\forall t((s\in s\wedge t\notin t)\rightarrow(s\neq t\wedge t\notin s))$.
+4. Axiom ‘specification’
+If $\Phi$ is definite and if exists set $u\in s$ such that $\Phi$ then it is not the case that for every set $u$ $(u\in s\rightarrow\neg\Phi)$.
+Thus we can postulate a set $v$ whose constituents are all and only those constituents of $s$ such that for every constituent of $v$ it is the case $\Phi$.
+That is axiom specification (also known as ‘separation’).
+  • if $u$ is free in $\Phi$ and $\Phi$ is definite then
+    $\forall s(\exists u(u\in s\wedge\Phi)\rightarrow\exists v\forall u(u\in v\leftrightarrow(u\in s\wedge\Phi)))$.
+Thus the conjunction of mutually exclusive properties does not constitute a set.
+Theorem.
+If an individual is a constituent of set $s$ then exists set $v$ whose
+constituents are all and only those individuals that are constituent of $s$.
+Proof. Formula $(u\in s\wedge u\in u)$ is definite and is not the conjunction of mutually
+exclusive properties. Apply specification to set $s$ to find set $v$
+$\exists v\forall u(u\in v\leftrightarrow(u\in s\wedge u\in u))$. Then an individual is a constituent of $v$.
+The formula applies to every $u$ thus we can choose $v$
+$\exists v(v\in v\leftrightarrow(v\in s\wedge v\in v))$ therefore
+$\exists v((v\notin v\vee(v\in s\wedge v\in v))\wedge((v\notin s\vee v\notin v)\vee v\in v))$.
+Then $\exists v(v\notin v\vee v\in s)$.
+Therefore $v$ is not an individual, or $v$ is an individual and a constituent of $s$. $\Box$
+Theorem.
+The set of all sets is nonexisting.
+Proof. If $\exists u(u\in s\wedge u\notin u)$ then formula $(u\in s\wedge u\notin u)$ is definite and is not the
+conjuction of mutually exclusive properties.
+Apply specification to set $s$ to find set $v$
+$\exists v\forall u(u\in v\leftrightarrow(u\in s\wedge u\notin u))$. Then a set is a constituent of $v$.
+The formula applies to every $u$ thus we can choose $v$
+$\exists v(v\in v\leftrightarrow(v\in s\wedge v\notin v))$ therefore
+$\exists v((v\notin v\vee(v\in s\wedge v\notin v))\wedge((v\notin s\vee v\in v)\vee v\in v))$.
+Then $\exists v(v\notin v\wedge v\notin s)$.
+Thus $v$ is not an individual and $v$ is not a constituent of $s$.
+Therefore the set of all sets is nonexisting. $\Box$
 5. Axiom ‘pairs’
@@
   • (pairs) $\forall s\forall t\exists v\forall u(u\in v\leftrightarrow(u=s\vee u=t))$.
-Set $v$ is denoted by $\{s,t\}$ thus $\{s,t\}=\{t,s\}$. The ‘singleton’ $\{s\}$ of $s$ is $\{s,s\}$.
+Set $v$ is denoted by $\{s,t\}$, $s$ and $t$ are the only constituents of $\{s,t\}$.
+Then $\{s,t\}=\{t,s\}$ thus $\{s,t\}$ is also called ‘an unordered pair’.
+The ‘singleton’ $\{s\}$ of set $s$ is $\{s,s\}$. Then
 $s\notin s\rightarrow\{s\}\neq s$ otherwise $s\in s$. Therefore $s\notin s\rightarrow\{s\}\notin\{s\}$,
@@
 However, if $s\neq t$ and $(s\in t\wedge t\in s)$ then $s\notin s$ and $t\notin t$.
-Then exists set $\{s,t\}$ thus there is the ‘vicious membership circle’ $(s\in t\wedge t\in s)$.
+Then exists set $\{s,t\}$ and the ‘vicious membership circle’ $(s\in t\wedge t\in s)$.
 That is contradictory to ‘the naming of individuals is logically first’.
-8. Axiom ‘regularity’
-The naming of individuals is logically first. Thus every set has a constituent, and vicious membership circles do not exist. Therefore we postulate ‘regularity’.
-  • (regularity) $\forall s\exists v(v\in s\wedge\forall u((u\in s\wedge u\in v)\rightarrow u\in u))$.
-If $s$ is an individual or a set of individuals then $v$ is an individual. Otherwise exists $v$ and $v$ is not an individual and $s$ and $v$ only have individuals in common.
-Then it is not the case that exists $s$ such that
-$\forall v(v\in s\rightarrow\exists u((u\in s\wedge u\in v)\wedge u\notin u))$.
-Therefore any sequence of sets $s$, $t$, $u$, $v$, … , finite (a vicious membership circle) or not finite, such that each term is a member of the previous one
-$t\in s$, $u\in t$, $v\in u$ is contradictory to ‘the naming of individuals is logically first’. Then $\forall s\forall t(s\neq t\rightarrow(s\notin t\vee t\notin s))$.
+Therefore if $s\neq t$ it is not the case that
+$\forall v(v\in\{s,t\}\rightarrow\exists u((u\in\{s,t\}\wedge u\in v)\wedge u\notin u))$.
+The naming of individuals is logically first. Thus vicious membership circles do not exist. We prove ‘regularity’.
+Theorem.
+It is not the case that exists set $w$ such that
+$\forall v(v\in w\rightarrow\exists u((u\in w\wedge u\in v)\wedge u\notin u))$.
+Proof. A finite sequence of sets $s$, $t$, $u$, …, $z$ such that $s\in t$, $t\in u$, …, $z\in s$
+is nonexisting otherwise there are vicious membership circles. And an ‘infinite’
+sequence of sets $s$, $t$, $u$, $v$, … such that each term is a member of the previous
+one $t\in s$, $u\in t$, $v\in u$ is nonexisting otherwise for every set $x$ belonging to
+the sequence exists set $y$ belonging to that sequence such that $y\in x$.
+That is contradictory to ‘the naming of individuals is logically first’. $\Box$
+Formula $\neg\exists w\forall v(v\in w\rightarrow\exists u((u\in w\wedge u\in v)\wedge u\notin u))$ is definite. Therefore
+  • (regularity) $\forall w\exists v(v\in w\wedge\forall u((u\in w\wedge u\in v)\rightarrow u\in u))$.
+The formula applies to every $u$ thus we can choose $v$. Then the formula reads
+$\forall w\exists v(v\in w\wedge((v\notin w\vee v\notin v)\vee v\in v))$, that is $\forall w\exists v(v\in w)$.
+Therefore an ‘empty set’ is nonexisting. If an individual is a constituent of $w$ then $v$ is an individual. Otherwise $v$ is not an individual and $w$ and $v$ only have individuals in common.
 6. Axiom ‘union’
@@
 If $s$ and $t$ are individuals then $s\cup t=\{s,t\}$.
-We prove: Any sequence of different sets $s$, $\bigcup s$, $\bigcup(\bigcup s)$, … is finite.
-Then $s\neq\bigcup s\rightarrow\exists t(t\in s\wedge t\notin t)$ otherwise $s=\bigcup s$.
+Theorem.
+Any sequence of different sets $s$, $\bigcup s$, $\bigcup(\bigcup s)$, … is finite.
+Proof. Then $s\neq\bigcup s\rightarrow\exists t(t\in s\wedge t\notin t)$ otherwise $s=\bigcup s$.
 $\bigcup s\neq\bigcup(\bigcup s)\rightarrow\exists u\exists t(u\in t\wedge t\in s\wedge u\notin u\wedge t\notin t)$ otherwise $\bigcup s=\bigcup(\bigcup s)$,
@@
 a member of the previous one $t\in s$, $u\in t$, $v\in u$, … contradictory to regularity.
-Therefore the sequence terminates thus exists $w$ belonging to the sequence and $\bigcup w=w$.
-If $z=s\cup\bigcup s\cup\bigcup(\bigcup s)\cup\ldots$ $w$ then $z$ is a set and $\bigcup z=z$.
-Set $z$ is called ‘the transitive closure’ of $s$.
-For every set the transitive closure exists.
-$\Box$
-The definition of equality and axiom pairs are only meaningful if exist sets the naming of which are logically first.
-Thus with individuals and axiom regularity.
+Therefore the sequence terminates. Thus exists set $w$ belonging to the sequence
+such that $\bigcup w=w$. If $z=s\cup\bigcup s\cup\bigcup(\bigcup s)\cup\ldots$ $w$ then $z$ is a set and $\bigcup z=z$.
+Set $z$ is called ‘the transitive closure’ of $s$. $\Box$
+The definitions of equality and axiom pairs are only meaningful if exist sets the naming of which are logically first.
+Thus with individuals and regularity.
 And regularity is used with axiom union to prove that for every set the transitive closure exists.
@@
 In (Halmos [1974] p.47) the Peano Axioms for the set of all natural numbers ”used to be considered as the fountainhead of all mathematical knowledge”.
-In this section we define ‘number sequences’ satisfying the Peano Axioms for natural numbers.
+In this section we define ‘transitive sets’ and ‘number sequences’ satisfying the Peano Axioms for natural numbers.
 Definition.
 Set $s$ is a ‘subset’ of $t$, denoted by $s\subseteq t$, if $\forall u(u\in s\rightarrow u\in t)$.
-$s$ is ‘transitive’ if $\forall u(u\in s\rightarrow u\subseteq s)$, or equivalently $\bigcup s\subseteq s$. $\Box$
+$s$ is transitive if $\forall u(u\in s\rightarrow u\subseteq s)$, or equivalently $\bigcup s\subseteq s$. $\Box$
 Every individual and every set of individuals is transitive. If $s$ is transitive then $\bigcup s$ and $s\cup\{s\}$ are transitive.
@@
 It is straightforward to define a number sequence satisfying the Peano Axioms for natural numbers.
-The first axiom of the Peano Axioms reads ‘there is a first natural number’ (denoted in (Mendelson [2015] p.153) by ‘00’).
+The first axiom of the Peano Axioms reads ‘there is a first natural number’.
+Natural numbers are also known as ‘finite ordinals’.
 Definition.
@@
 Therefore ‘the set of all natural numbers’ is nonexisting.
-However, to postulate ‘set $\omega$ of all natural numbers with first number $\alpha$’ is not excluded.
+However, it is not excluded to postulate ‘set $\omega$ of all natural numbers with first number $\alpha$’.
 Then $\bigcup\omega=\omega$.
-Therefore we can define ‘transfinite ordinal numbers with first number $\omega$’ analogous to natural numbers (the finite ordinal numbers) with first number $\alpha$.
-In the Introduction (Wittgenstein [1999] p.21) Bertrand Russell writes
-”No logic can be considered adequate until it has been shown to be capable of dealing with transfinite numbers”.
-The logic in Very basic set theory is capable to deal with transfinite numbers.
+Therefore we can define ‘transfinite ordinal numbers with first number $\omega$’ analogous to natural numbers with first number $\alpha$.
+Therefore the logic in Very basic set theory is capable to deal with transfinite numbers.
 Conclusion
@@
 It is a set theory with individuals, the logic is capable to deal with transfinite numbers and the naming of individuals is logically first.
-But individuals are not needed for mathematical purposes therefore individuals are banished from ‘mainstream set theory’.
+There is no ‘empty set’ and regularity is not an axiom but a theorem.
 There are more differences with mainstream set theory. For example, the axiom ‘powerset’ is as usual but Cantor’s Theorem does not hold.

### v8

@@
 $\forall w\exists v(v\in w\wedge((v\notin w\vee v\notin v)\vee v\in v))$, that is $\forall w\exists v(v\in w)$.
-Therefore an ‘empty set’ is nonexisting. If an individual is a constituent of $w$ then $v$ is an individual. Otherwise $v$ is not an individual and $w$ and $v$ only have individuals in common.
+Therefore an ‘empty set’ is nonexisting.
+If an individual is a constituent of $w$ then $v$ is an individual.
+And if $\exists v(v\notin v\wedge v\in w)$ then $\exists v(v\notin v\wedge v\in w)$ such that $v$ and $w$ only have individuals in common.
 6. Axiom ‘union’
-The members of the members of a set constitute a set. We postulate ‘union’.
+The constituents of the members of a set constitute a set. We postulate ‘union’.
   • (union) $\forall s\exists v\forall u(u\in v\leftrightarrow\exists t(t\in s\wedge u\in t))$.
@@
 $\bigcup(\bigcup s)\neq\bigcup(\bigcup(\bigcup s))\rightarrow\exists v\exists u\exists t(v\in u\wedge u\in t\wedge t\in s\wedge v\notin v\wedge u\notin u\wedge t\notin t)$,
-Then there is an ‘infinite’ sequence of sets $s$, $t$, $u$, $v$, … such that each term is
-a member of the previous one $t\in s$, $u\in t$, $v\in u$, … contradictory to regularity.
+Thus there is a rule to construct for every set $z\neq s$ belonging to the sequence a
+series of sets $t$, $u$, $v$, , $x$, $y$, $z$ such that each term is a member of the previous
+one: $t\in s$, $u\in t$, $v\in u$, , $x\in v$, $y\in x$, $z\in y$ contradictory to regularity if
+the sequence is ‘infinite’: the naming of individuals is logically first.
 Therefore the sequence terminates. Thus exists set $w$ belonging to the sequence
-such that $\bigcup w=w$. If $z=s\cup\bigcup s\cup\bigcup(\bigcup s)\cup\ldots$ $w$ then $z$ is a set and $\bigcup z=z$.
-Set $z$ is called ‘the transitive closure’ of $s$. $\Box$
+such that $\bigcup w=w$. Then $s\cup\bigcup s\cup\bigcup(\bigcup s)\cup\ldots$ $w$ is a set, and is called ‘the
+transitive closure of $s$’. $\Box$
 The definitions of equality and axiom pairs are only meaningful if exist sets the naming of which are logically first.
@@
 And regularity is used with axiom union to prove that for every set the transitive closure exists.
-The proof needs no other axioms and no ‘set of all natural numbers’.
+The proof requires no other axioms or no ‘set of all natural numbers’.
 7. Natural numbers
@@
 Therefore ‘the set of all natural numbers’ is nonexisting.
-However, it is not excluded to postulate ‘set $\omega$ of all natural numbers with first number $\alpha$’.
+However, it is not prohibited to postulate ‘set $\omega$ of all natural numbers with first number $\alpha$’.
 Then $\bigcup\omega=\omega$.
-Therefore we can define ‘transfinite ordinal numbers with first number $\omega$’ analogous to natural numbers with first number $\alpha$.
-Therefore the logic in Very basic set theory is capable to deal with transfinite numbers.
+Therefore we can define a number sequence with first number $\omega$ analogous to natural numbers with first number $\alpha$.
+Then there are the finite ordinals (natural numbers) with first number $\alpha$, followed by ‘transfinite ordinal numbers with first number $\omega$’.
+And if we define $\omega+\omega=\omega\cdot 2$ then $\bigcup(\omega\cdot 2)=\omega\cdot 2$ thus the next sequence of transfinite ordinals are the ordinals with first number $\omega\cdot 2$, and so on, that is
+$\alpha$, $\alpha+1$, $\alpha+2$, …, $\omega$, $\omega+1$, $\omega+2$, …, $\omega\cdot 2$, $\omega\cdot 2+1$, $\omega\cdot 2+2$, …, $\omega\cdot 3$, ….
+Thus the logic in Very basic set theory is capable to deal with transfinite numbers.
 Conclusion
-Very basic set theory is a set theory that has its philosophical foundation in Ludwig Wittgenstein’s Tractatus Logico-Philosophicus.
-It is a set theory with individuals, the logic is capable to deal with transfinite numbers and the naming of individuals is logically first.
-There is no ‘empty set’ and regularity is not an axiom but a theorem.
-There are more differences with mainstream set theory. For example, the axiom ‘powerset’ is as usual but Cantor’s Theorem does not hold.
-  • (powerset) $\forall s\exists v\forall u(u\in v\leftrightarrow u\subseteq s)$.
-Set $v$ is denoted by ${\mathcal{P}}s$. Then ${\mathcal{P}}\{s\}=\{\{s\}\}$ thus the ‘number of members’ of $\{s\}$
-is equal to the ‘number of members’ of ${\mathcal{P}}\{s\}$.
+Very basic set theory is a set theory with the philosophical foundation in Ludwig Wittgenstein’s Tractatus Logico-Philosophicus.
+It is a set theory with individuals.
+The atomic facts are the membership relations and the elementary propositions are the individuals.
+The formulas of set theory preserve definiteness.
+The naming of individuals is logically first and the logic is capable to deal with transfinite numbers.
+We list some differences with ‘mainstream set theory’.
+  • There is no ‘empty set’ (Kanamori [2003] §1).
+  • The singleton of an individual is that individual itself (Kanamori [2003] §2).
+  • Regularity is not an axiom but a theorem.
+  • The proof that ‘every set has a transitive closure’ requires no other axioms or no ‘set of all natural numbers’ (Mendelson [2015] pp.286-287).
   • Cantor’s Theorem (Zermelo [1908] p.276) reads
     for every set $s$, the ‘number of members’ of $s$ is less than the ‘number of
-members’ of ${\mathcal{P}}s$.
-Thus Cantor’s Theorem does not hold. $\Box$
+    members’ of the powerset ${\mathcal{P}}s$.
+    However, there is no ‘empty set’ therefore ${\mathcal{P}}\{s\}=\{\{s\}\}$.
+    Thus Cantor’s Theorem does not hold.
+  • Any pair of individuals is the first number of a sequence of natural numbers.
+Then transfinite ordinals are postulated.
+If e.g.
+$\bigcup\omega=\omega$ then $\omega$ is the first number of a sequence of transfinite ordinals.
+Therefore $\omega$ is not a ‘limit ordinal’, $\omega$ is the first number of a sequence satifying the Peano Axioms for natural numbers.